### v1

# The resolution quiver of Nakayama algebras which are minimal Auslander-Gorenstein

Abstract.
Let $A$ be a Nakayama algebra. Using Ringel’s resolution quiver, we give a criterion to decide whether $A$ is minimal Auslander-Gorenstein. The criterion strongly relies on the parity of the selfinjective dimension of $A$.
Key words and phrases:
Nakayama algebra, Auslander-Gorenstein algebra, resolution quiver
2020 Mathematics Subject Classification:
Primary 16E10; Secondary 16G20
1. Introduction
Following Iyama [4], an algebra is said to be higher Auslander if its global dimension is finite and bounded by its dominant dimension. Higher Auslander algebras arise in cluster-tilting theory. Following Iyama and Solberg [5], an algebra is said to be minimal Auslander-Gorenstein provided that its global dimension is finite and bounded by its dominant dimension. Minimal Auslander-Gorenstein algebras are a special class of Auslander-Gorenstein algebras.
Nakayama algebras are one of most basic classes of algebras in representation theory of algebras. An algebra is called a Nakayama algebra if it is a homomorphic image of the path algebra of an oriented cycle. A connected Nakayama algebras is either linear or else cyclic.
The resolution quiver for Nakayama algebras is introduced by Ringel [10]. It is an efficient tool for studying the homological properties of Nakayama algebras. A resolution quiver criterion to detect the finiteness of global dimension and the finiteness of selfinjective dimension for Nakayama algebras is given in [16].
The syzygy filtered algebras for cyclic Nakayama algebras is introduced by Sen [12]. It plays a role in constructing cyclic Nakayama algebras which are higher Auslander in [13]. The reversed syzygy filtered algebras for all Nakayama algebras is introduced in [15]. It provides a systematic construction of cyclic Nakayama algebras which are minimal Auslander-Gorenstein and higher Auslander.
Inspired by their work, we study syzygy filtered algebras for all Nakayama algebras. Then we give a resolution quiver criterion to decide whether a Nakayama algebra is minimal Auslander-Gorenstein. This criterion strongly depends on the parity of the selfinjective dimension.
Let $A$ be a Nakayama algebra with translation $\tau$. For an $A$-module $M$, denote by $P(M)$ the projective cover and by $I(M)$ the injective envelope of $M$. The resolution quiver $R(A)$ of $A$ is defined as follows. The vertex set of $R(A)$ is the isoclasses of all simple $A$-modules, and there is a unique arrow from $S$ to $\tau\operatorname{soc}P(S)$ for each simple $A$-module $S$. A simple $A$-module $S$ is said to be black if the projective dimension of $S$ is not equal to $1$, and $S$ is said to be red otherwise. This assigns colors to the vertices of resolution quivers.
The weight of a cycle $C$ in $R(A)$ is the sum of $c_{a}/n$, where $a$ runs through all vertices in $C$ and $n$ is rank of $A$. The weight is an integer and all cycles in $R(A)$ have the same weight. The weight of any cycle in $R(A)$ is called the weight of $R(A)$.
A simple $A$-module $S$ said to be cyclic if $S$ lies on cycle in $R(A)$. For any simple $A$-module $S$, denote by $\operatorname{dist}S$ the minimal integer $d$ such that the $d$-th successor of $S$ is cyclic in $R(A)$. Let $\operatorname{dist}A$ be the maximum of the numbers $\operatorname{dist}S$, where $S$ runs through all simple $A$-modules.
The following is the main result of this paper.
Theorem A.
Let $A$ be a Nakayama algebra and let $R(A)$ be the resolution quiver of $A$. Then $A$ is minimal Auslander-Gorenstein of odd selfinjective dimension if and only if
  1. (1)
    $\operatorname{dist}S=\operatorname{dist}A$ for every leaf $S$ in $R(A)$;
  2. (2)
    every noncyclic vertex has at most $1$ predecessor in $R(A)$;
  3. (3)
    $R(A)$ is connected of weight $1$;
  4. (4)
    $\tau^{-1}$ of all cyclic vertices in $R(A)$ are black.
Theorem B.
Let $A$ be a Nakayama algebra and let $R(A)$ be the resolution quiver of $A$. Then $A$ is minimal Auslander-Gorenstein of even selfinjective dimension if and only if
  1. (1)
    $\operatorname{dist}S=\operatorname{dist}A$ for every leaf $S$ in $R(A)$;
  2. (2)
    every noncyclic vertex has at most $1$ predecessor in $R(A)$;
  3. (3)
    every cyclic vertex has at most $2$ predecessor in $R(A)$;
  4. (4)
    all cyclic vertices in $R(A)$ are black.
Throughout this paper, $k$ is a fixed algebraically closed field. All algebras are finite dimensional associative algebras over $k$. All modules are finite dimensional left modules unless otherwise stated.
The paper is organized as follows. In Section 2, we recall some basic properties of Nakayama algebras. In Section 3, we study syzygy filtered algebra for all Nakayama algebras. Then we prove the criterion in Section 4. Some applications and examples is given in Section 5.
Acknowledgements
The author is grateful to Professor Shijie Zhu for his comments and suggestions. This work is supported by the National Natural Science Foundation of China (No. 1257011205).
2. Preliminary
We refer to [1] for quivers and their representations.
Given an integer $n$, we denote by $\Delta_{n}$ the quiver on $n$ vertices containing a single directed cycle through all vertices; see Figure 1. Denote by $k\Delta_{n}$ the path algebra of the quiver $\Delta_{n}$.
112233nn
Figure 1. The quiver $\Delta_{n}$.
Following [3], an algebra $A$ is said to be a Nakayama algebra of rank $n$ if it is isomorphic to the quotient algebra of $k\Delta_{n}$ via the ideal generated by a finite set $I$ of nontrivial paths. If $I$ consists of paths of length at least $2$, then $A$ is a cyclic Nakayama algebra. If there exists a unique arrow in $I$, then $A$ is a linear Nakayama algebra. Both cyclic and linear Nakayama algebras are connected. We allow more than one arrow in $I$.
Let $A$ be a Nakayama algebra of rank $n$. For each vertex $a$ in $\Delta_{n}$, denote by $S(a)$ the simple $A$-module, $P(a)$ the projective $A$-module and $I(a)$ the injective $A$-module at $a$. Set $\tau S(a)=S(b)$ if there is an arrow from $a$ to $b$ in $\Delta_{n}$. The map $\tau$ yields a translation of simple $A$-modules. If $A$ is a cyclic Nakayama algebra, then $\tau$ coincides with the Auslander-Reiten translation.
For each vertex $a$ in $\Delta_{n}$, denote by $c_{a}$ the composition length of $P(a)$. The series
$c(A)=(c_{a},c_{\tau a},\cdots,c_{\tau^{n-1}a})$
is called a Kupisch series of $A$. Two Nakayama algebras $A$ and $B$ are isomorphic if and only if $c(A)$ and $c(B)$ are equal up to a power of cyclic permutation.
Denote by $R(A)$ the resolution quiver of $A$. The vertices of $R(A)$ are the isoclasses of all simple $A$-modules, and there is a unique arrow from $S$ to
$\gamma(S)=\tau\operatorname{soc}P(S)$
for each simple $A$-module $S$; see also [2]. A simple $A$-module $S$ is said to be black if the projective dimension of $S$ is not equal to $1$, and $S$ is said to be red otherwise. This assigns colors to the vertices of $R(A)$.
The following is well known.
Lemma 2.1.
Let $S$ be a simple $A$-module. Then
  1. (1)
    $P(S)$ is injective if and only if $\tau^{-1}S$ is black;
  2. (2)
    $\gamma(\tau S)=\gamma(S)$, provided that $S$ is red.
Corollary 2.2.
Let $S$ be a simple $A$-module. Then $S$ and $\tau^{-1}S$ are black if and only if $\gamma^{-1}(\gamma S)=\{S\}$.
We need the following; see [16, 1.1] and [11, 3.10].
Lemma 2.3.
Let $A$ be a Nakayama algebra. Then
  1. (1)
    AA has finite global dimension if and only if $R(A)$ is connected of weight $1$;
  2. (2)
    $2\operatorname{dist}A-1\leq\mathrm{gl}.\dim A\leq 2\operatorname{dist}A$, provided that $\mathrm{gl}.\dim A$ is finite.
An algebra $A$ is said to be Gorenstein if the left selfinjective dimension and right selfinjective dimension of $A$ are finite. For any Gorenstein algebra, two selfinjective dimensions coincide; see [17]. It is known that a Nakayama algebra is Gorenstein if and only if its left selfinjective dimension is finite.
We need the following; see [16, 1.2] and [12, 4.13].
Lemma 2.4.
Let $A$ be a Nakayama algebra of infinite global dimension. Then
  1. (1)
    AA is Gorenstein if and only if all cycles in $R(A)$ are black;
  2. (2)
    $\mathrm{id}_{A}\,A=2\operatorname{dist}A$, provided that $\mathrm{id}_{A}\,A$ is finite.
Let $A$ be a Nakayama algebra and $R(A)$ be its resolution quiver. A vertex $S$ in $R(A)$ is said to be a leaf if there is no arrow ending with $S$.
We need the following; see [6, 3.1].
Lemma 2.5.
Let $S$ be a simple $A$-module. Then $\mathrm{id}_{A}\,S=1$ if and only if $S$ is a leaf in $R(A)$.
3. Syzygy filtered algebra
In the present section, we study the syzygy filtered algebra for all Nakayama algebras and give some necessary results for the proof of the main result.
Let $A$ be a Nakayama algebra of rank $n$ and let $R(A)$ be the resolution quiver of $A$. Following [11, A.6] the map $\gamma$ induces a one-to-one correspondence between the set of all black vertices and the set of all non-leaf vertices in $R(A)$. For any non-leaf vertex $S$ in $R(A)$, let $T$ be the unique black vertex such that $\gamma(T)=S$. Consider the exact sequence
$0\to\Delta(S)\to P(\tau T)\to P(T)\to T\to 0.$ (3.1)
The base module $\Delta(S)$ is the quotient module of $P(S)$ of minimal length such that the translation of its socle is a non-leaf vertex.
For any module $M$, denote by $\Omega(M)$ the syzygy module of $M$. Then we have
$\Delta(S)=\begin{cases}P(S),&\text{if}\;\mathrm{pd}_{A}\,T\in\{0,2\},\\ \Omega^{2}(T),&\text{otherwise}.\end{cases}$ (3.2)
Following [12, 14], the base set
$\mathcal{B}=\{\Delta(S)\mid\text{$S$ is a non-leaf vertex in $R(A)$}\}$
form a semibrick in the category $\operatorname{mod}A$ of all finite dimensional $A$-modules. The filtration category $\mathcal{F}$ of $\mathcal{B}$ is an exact Abelian subcategory of $\operatorname{mod}A$. The category $\mathcal{F}$ contains the projective cover $\mathcal{P}$ of all base modules and the second syzygy of all finite dimensional $A$-modules.
The syzygy filtered algebra $\varepsilon(A)$ is the endomorphism algebra of $\mathcal{P}$, that is,
$\varepsilon(A)=\operatorname{End}_{A}(\bigoplus\nolimits_{P\in\mathcal{P}}P).$
It is a Nakayama algebra of rank at most $n$. The equality holds if and only if $A$ is selfinjective. The category $\operatorname{mod}\varepsilon(A)$ is equivalent to $\mathcal{F}$. The algebra $\varepsilon(A)$ has finite global dimension if and only if $A$ has finite global dimension.
For a Nakayama algebra $A$, denote by $\chi(A)$ the number of isoclasses of simple projective $A$-modules. By (3.2), we have the following.
Lemma 3.1.
Let $A$ be a Nakayama algebra. Then $\chi(\varepsilon{A})-\chi(A)$ is equal to the number of isoclasses of simple $A$-modules of projective dimension $2$.
For any integer $i\geq 0$, denote by $\varepsilon^{i}(A)$ the $i$-th iterated syzygy filtered algebra of $A$. We have the following.
Corollary 3.2.
For any $i\geq 1$, $\chi(\varepsilon^{i}{A})-\chi(\varepsilon^{i-1}A)$ is equal to the number of isoclasses of simple $A$-modules of projective dimension $2i$.
There is an integer $r$ such that $\varepsilon^{r}(A)$ is a selfinjective Nakayama algebra. Then $\chi(\varepsilon^{r}{A})$ is equal to the number of isoclasses of simple $A$-modules of even projective dimension. This gives an alternative proof of Madsen’s criterion [6], which states that a Nakayama algebra has finite global dimension if and only if it has a simple module of even projective dimension.
The projective objects in $\mathcal{F}$ are precisely the projective $A$-modules which are contained in $\mathcal{F}$. However, the injective objects in $\mathcal{F}$ are more complicated.
Following [7], the defect $\operatorname{def}A$ of an algebra $A$ is the number of isoclasses of indecomposable injective non-projective $A$-modules. If $A$ is a Nakayama algebra, then $\operatorname{def}A$ is equal to the number of red vertices in its resolution quiver. By (3.2) we have following.
Lemma 3.3.
For any $i\geq 0$, $\operatorname{def}\varepsilon^{i}(A)$ is equal to the number of isoclasses of simple $A$-modules of projective dimension $2i+1$.
Let $S$ be a non-leaf vertex in $R(A)$. By [11, A.6] there is a unique vertex $T$ such that $P(T)$ is injective and $\gamma(T)=S$. Consider the exact sequence
$0\to J(\tau^{-1}S)\to P(T)\to P(I(\tau^{-1}T))\overset{f}{\to}I(\tau^{-1}T)\to 0.$ (3.3)
We have the following; compare [12, 3.18].
Lemma 3.4.
The assignment $S\mapsto J(\tau^{-1}S)$ gives a bijection between the isoclasses of simple $A$-modules of injective dimension not equal to $1$ and the isoclasses of indecomposable injective objects in $\mathcal{F}$. It restricts to a bijection between the isoclasses of simple $A$-modules of injective dimension $3$ and the isoclasses of indecomposable injective non-projective objects in $\mathcal{F}$.
Proof.
If $T$ is a non-leaf vertex, the module $I(\tau^{-1}T)$ is projective. The sequence (3.3) splits and $J(\tau^{-1}S)$ is an indecomposable injective $A$-module contained in $\mathcal{F}$.
If $T$ is a leaf, then $I(\tau^{-1}T)$ is not projective and $\operatorname{Ker}f$ is nonzero. Since $P(T)$ is an injective $A$-module, $J(\tau^{-1}S)$ is nonzero. Then $J(\tau^{-1}S)$ is an indecomposable object in $\mathcal{F}$. Since $\operatorname{Ker}f$ is a proper submodule of a base module and $P(T)$ is an injective $A$-module, we know that $J(\tau^{-1}S)$ is an injective object in $\mathcal{F}$.
The number of isoclasses of injective objects in $\mathcal{F}$ is equal to the number of non-leaf vertices in $R(A)$. If $S$ and $S^{\prime}$ are not isomorphic, $J(\tau^{-1}S)$ and $J(\tau^{-1}S^{\prime})$ are not isomorphic. Then the first conclusion holds.
Note that $J(\tau^{-1}S)$ is a non-projective object in $\mathcal{F}$ if and only if $\Delta(S)$ is a leaf in $R(\varepsilon{A})$. This happens if and only if the injective dimension of $S$ is $3$. ∎
The translation $\tau$ induces a translation $\tau_{\varepsilon}$ of simple objects in $\mathcal{F}$. For any vertex $\Delta(S)$ in $R(\varepsilon{A})$, by [11, B.4] we have
$\gamma_{\varepsilon}(\Delta(S))=\tau_{\varepsilon}\operatorname{soc}_{\mathcal{F}}P(\Delta(S))=\Delta(\gamma S).$
The assignment $\Delta(S)\mapsto S$ yields an injective map
$\varphi\colon R(\varepsilon{A})\to R(A)$
of resolution quivers. Denote by $R(A)^{\prime}$ the image of $\varphi$. We mention that $\varphi$ preserves the weight of any cycle. However, $\varphi$ may not preserve the vertex colors.
A vertex $T$ in $R(A)$ is said to be *-black if $\tau^{-1}T$ is black, and $T$ is said to be *-red if $\tau^{-1}T$ is red. For each leaf $S$ in $R(A)^{\prime}$, there is a unique black leaf $T_{0}$ in $R(A)$ such that $\gamma(T_{0})=S$ and a unique *-black leaf $T_{1}$ in $R(A)$ such that $\gamma(T_{1})=S$. For any resolution quiver $R$, denote by $\mathrm{Leaf}\mbox{-}R,\mathrm{BLeaf}\mbox{-}R$ and $\mathrm{*BLeaf}\mbox{-}R$ the set of all leaf vertices, black leaf vertices and *-black leaf vertices, respectively. There is an injective map
$\varphi_{0}\colon\mathrm{Leaf}\mbox{-}R(A)^{\prime}\to\mathrm{BLeaf}\mbox{-}R(A)$
sending $S$ to $T_{0}$ and an injective map
$\varphi_{1}\colon\mathrm{Leaf}\mbox{-}R(A)^{\prime}\to\mathrm{*BLeaf}\mbox{-}R(A)$
sending $S$ to $T_{1}$. Consider the following two conditions.
  1. (B)
    For each black leaf in $R(A)$, each predecessor of its successor is a leaf.
  2. (B)
    For each *-black leaf in $R(A)$, each predecessor of its successor is a leaf.
Then $A$ satisfies (B) if and only if the map $\varphi_{0}$ is bijective, and $A$ satisfies (B) if and only if the map $\varphi_{1}$ is bijective.
We have the following.
Proposition 3.5.
Let $A$ be a Nakayama algebra. The following are equivalent.
  1. (1)
    AA satisfies (B);
  2. (2)
    $\varphi$ preserves vertex colors;
  3. (3)
    The number of red vertices in $R(\varepsilon{A})$ and that $R(A)^{\prime}$ are equal.
Proof.
$(1)\iff(3)$ Recall that the number of red vertices is equal to the number of leaf vertices in any resolution quiver. We have
$|\mathrm{Red}\mbox{-}R(A)^{\prime}|=|\mathrm{Red}\mbox{-}R(A)|-|\mathrm{RLeaf}\mbox{-}R(A)|=|\mathrm{BLeaf}\mbox{-}R(A)|$
and
$|\mathrm{Red}\mbox{-}R(\varepsilon{A})|=|\mathrm{Leaf}\mbox{-}R(\varepsilon{A})|=|\mathrm{Leaf}\mbox{-}R(A)^{\prime}|\leq|\mathrm{BLeaf}\mbox{-}R(A)|.$
The equality holds if and only if $\varphi_{0}$ is a bijection.
$(2)\iff(3)$ Let $P$ be an indecomposable projective object in $\mathcal{F}$. If $P$ is minimal in $\operatorname{mod}A$, so is $P$ in $\mathcal{F}$. This shows that $\varphi$ preserves red vertices. Then $\varphi$ preserves vertex colors if and only if the numbers of red vertices in $R(\varepsilon{A})$ and $R(A)^{\prime}$ are equal. ∎
Corollary 3.6.
Let $A$ be a Nakayama algebra satisfying (B). Then $A$ is Gorenstein if and only if $\varepsilon(A)$ is Gorenstein. Moreover, we have
$\mathrm{id}_{A}\,A=\mathrm{id}_{\varepsilon(A)}\,\varepsilon(A)+2$ (3.4)
provided that $A$ is not selfinjective.
Proof.
Since $A$ satisfies (B), then $\varphi$ preserves vertex colors. Recall that $A$ has finite global dimension if and only if $\varepsilon(A)$ has finite global dimension. Then by Lemma 2.4 the algebra $A$ is Gorenstein if and only if $\varepsilon(A)$ is Gorenstein. Note that the selfinjective dimension of $A$ is not $1$. The equation (3.4) follows from [12, 3.23]. ∎
Example 3.7.
Let $A$ be a Nakayama algebra with Kupisch series $(3,2,2)$. The resolution quiver $R(A)$ is the following. One checks that $A$ satisfies (B).
123
Note that $\Delta(1)=S(1)$ and $\Delta(2)=P(2)$. The Kupisch series of $\varepsilon(A)$ is $(2,1)$. The resolution quiver of $\varepsilon(A)$ is the following.
12
The colors of $1$ and $2$ are preserved.
Example 3.8.
Let $A$ be a Nakayama algebra with Kupisch series $(3,2,1)$. The resolution quiver $R(A)$ is the following. One sees that $A$ does not satisfy (B).
123
We have $\Delta(1)=P(1)$. The Kupisch series of $\varepsilon(A)$ is $(1)$. The resolution quiver of $\varepsilon(A)$ is the following.
1
The color of $1$ is not preserved.
Similarly, we have the following.
Proposition 3.9.
Let $A$ be a Nakayama algebra. The following are equivalent.
  1. (1)
    AA satisfies (B);
  2. (2)
    $\varphi$ preserves vertex *-colors;
  3. (3)
    The number of *-red vertices in $R(\varepsilon{A})$ and $R(A)^{\prime}$ are equal.
Example 3.10.
Let $A$ be a Nakayama algebra with Kupisch series $(8,7,7,6)$. The resolution quiver of $R(A)$ is the following. It is easy to see that $A$ satisfies (B).
1234
We have two exact sequences
$0\to J(4)\to P(1)\to P(3)\to I(1)\to 0,$
$0\to J(1)\to P(3)\to P(1)\to I(2)\to 0.$
Note that $J(4)$ is isomorphic to $P(1)$ and $J(1)$ is an indecomposable injective non-projective object in $\mathcal{F}$.
4. Proof of the main result
In this section, we study the dominant dimensions of Nakayama algebras. Then we prove the main result of the present paper.
Let $A$ be an algebra and $M$ be an $A$-module. Take a minimal injective resolution
$0\to M\to I^{0}\to I^{1}\to\cdots\to I^{n}\to\cdots$
of $M$. The dominant dimension $\mathrm{dom}.\dim_{A}M$ of $M$ is the maximal integer $n$ or $\infty$ such that $I^{j}$ is projective for all $j<n$. Dually, take a minimal projective resolution
$\cdots\to P_{n}\to\cdots P_{1}\to P_{0}\to M\to 0$
of $M$. The codominant dimension $\mathrm{codom}.\dim_{A}M$ of $M$ is the maximal integer $n$ or $\infty$ such that $P_{j}$ is injective for all $j<n$. The left dominant dimension of $A$ is equal to the right dominant dimension of $A$; see [8]. It is known that any Nakayama algebra has dominant dimension at least $1$.
For an integer $d\geq 1$, an algebra $A$ is called higher Auslander if
$\mathrm{dom}.\dim A=d=\mathrm{gl}.\dim A,$
and $A$ is called minimal Auslander-Gorenstein if
$\mathrm{dom}.\dim A=d=\mathrm{id}_{A}\,A.$
For algebras of selfinjective dimension $1$, we have the following.
Proposition 4.1.
Let $A$ be a Nakayama algebra and $R(A)$ be its resolution quiver. Then $A$ is minimal Auslander-Gorenstein of selfinjective dimension $1$ if and only if
  1. (1)
    $\operatorname{dist}A=1$;
  2. (2)
    $R(A)$ is connected of weight $1$;
  3. (3)
    $\tau^{-1}$ of all cyclic vertices in $R(A)$ are black.
Proof.
$\implies$" Since $\mathrm{id}_{A}\,A=1$, we have $\mathrm{gl}.\dim A=1$ by Lemma 2.4. Then by Lemma 2.3 $R(A)$ is connected of weight $1$. There is a direct sum decomposition
$A=A_{1}\oplus A_{2}\oplus\cdots\oplus A_{r}$
where $A_{i}$ is a hereditary linear Nakayama algebra for all $i\leq r$. Let $S$ be a cyclic vertex in $R(A)$. Then $S$ is the simple module at the source of $A_{i}$ and $\tau^{-1}{S}$ is the simple module at the sink of $A_{i-1}$. Since $\tau^{-1}S$ is projective, it is black.
[…]
$0\to\Delta(S)\to P(\tau T)\to P(T)\to T\to 0$
such that $\tau^{-1}S$ is black. Then $P(S)$ is injective by Lemma 2.1. Since $\mathrm{gl}.\dim A$ is at most $2$, by (3.2) we have $\Delta(S)=P(S)$. Then the sequence splits and $T$ is projective. We conclude that the selfinjective dimension of $A$ is $1$. ∎
Remark 4.2.
Let $A$ be a Nakayama algebra of global dimension $1$. Since not all components of $A$ are simple, there is at least one red cyclic vertex in $R(A)$.
Example 4.3.
Let $A$ be a Nakayama algebra with Kupisch series $(1,3,2,1)$. Then $A$ is the direct sum of two hereditary linear Nakayama algebras. The resolution quiver of $R(A)$ is the following. Observe that the cyclic vertex $2$ is a red vertex.
1234
For algebras of dominant dimension at least $2$, we need the following.
Lemma 4.4.
Let $A$ be a Nakayama algebra of dominant dimension at least $2$ and let $S$ be a simple $A$-module. If $\mathrm{pd}_{A}\,\tau^{-1}S=1$, then $\mathrm{id}_{A}\,S\neq 1$.
Proof.
Since $\mathrm{pd}_{A}\,\tau^{-1}S=1$, then $\tau^{-1}S$ is red. Let $m$ be the minimal integer such that $\tau^{-m-1}S$ is black. By Lemma 2.1 we know that $\gamma S=\gamma\tau^{-m}S$ and $P(\tau^{-m}S)$ is injective. Consider the exact sequence
$0\to P(S)\to P(\tau^{-m}S)\to I(\tau^{-1}S).$
Since $\mathrm{dom}.\dim A\geq 2$, the module $I(\tau^{-1}S)$ is projective. Then the injective dimension of $S$ is not equal to $1$. ∎
We have the following.
Proposition 4.5.
Let $A$ be a Nakayama algebra and $R(A)$ be its resolution quiver. Then $A$ is minimal Auslander-Gorenstein of selfinjective dimension $2$ if and only if
  1. (1)
    $\operatorname{dist}A=1$;
  2. (2)
    $|\gamma^{-1}(S)|\leq 2$ for every cyclic vertex $S$ in $R(A)$;
  3. (3)
    all cyclic vertices in $R(A)$ are black.
Proof.
$\implies$” (1) If $A$ has finite global dimension, then $\operatorname{dist}A=1$ by Lemma 2.3. If $A$ has infinite global dimension, by Lemma 2.4 we have $\operatorname{dist}A=1$.
(2) Suppose three distinct vertices $T,\tau T,\tau^{2}T$ have the same cyclic successor in $R(A)$. Then $T$ and $\tau T$ are red. Since $\mathrm{dom}.\dim A\geq 2$, by Lemma 4.4 $\tau T$ and $\tau^{2}T$ are non-leaf vertices. By (1) we get that $\tau T,\tau^{2}T$ are cyclic. This is impossible.
(3) Let $S$ be a red vertex in $R(A)$. By (2) $\tau^{-1}S$ is black. By Lemma 2.1 $P(S)$ is injective. Since the selfinjective dimension of $A$ is $2$, there is an exact sequence
$0\to P(\tau S)\to P(S)\to I(S)\to I(\tau^{-1}S)\to 0.$
Then the injective dimension of $S$ is equal to $1$. By Lemma 2.5 the vertex $S$ is a leaf in $R(A)$. In particular, $S$ is not cyclic.
[…]
By Remark 4.2 the algebra $A$ is not hereditary. Then the selfinjective dimension of $A$ is $2$. It remains to prove $\mathrm{dom}.\dim A\geq 2$. By [9, 6.2], it suffices to show that each red simple $A$-module is in the image of black simple $A$-modules under $\tau^{-1}\gamma\tau$.
Since $\operatorname{dist}A=1$ and all cyclic vertices are black. We infer that all black vertices are cyclic. Let $S$ be a red vertex in $R(A)$. Since $\gamma(S)$ has at most two predecessors, $\tau(S)$ is black and then cyclic. There is a cyclic vertex $T$ in $R(A)$ such that $\gamma(T)=\tau(S)$. If $\tau^{-1}T$ is black, then
$\tau^{-1}\gamma\tau(\tau^{-1}T)=\tau^{-1}\gamma T=\tau^{-1}\tau S=S.$
If $\tau^{-1}T$ is red, then $\tau^{-2}T$ is black. We have $\gamma\tau^{-1}T=\gamma T$ and
$\tau^{-1}\gamma\tau(\tau^{-2}T)=\tau^{-1}\gamma\tau^{-1}T=\tau^{-1}\gamma T=\tau^{-1}\tau S=S.$
This finishes our proof. ∎
Example 4.6.
Let $A$ be a Nakayama algebra with Kupisch series $(6,7,6,7)$. Then $R(A)$ is the following.
1234
It is easy to see that $R(A)$ fulfills the condition in Proposition 4.5. Then $A$ is minimal Auslander-Gorenstein of selfinjective dimension $2$.
For algebras of dominant dimension at least $3$, we need the following key result; see [15, 4.1 and 4.2]. We give a short proof here.
Lemma 4.7.
Let $A$ be a Nakayama algebra. The following are equivalent.
  1. (1)
    $\mathrm{dom}.\dim A\geq 3$;
  2. (2)
    $\tau^{-1}S$ and $S$ are black for every leaf $S$ in $R(A)$;
  3. (3)
    $\gamma^{-1}(\gamma S)=\{S\}$ for every leaf $S$ in $R(A)$;
  4. (4)
    $\operatorname{def}A=\operatorname{def}\varepsilon(A)$;
  5. (5)
    $\mathrm{dom}.\dim A=\mathrm{dom}.\dim\varepsilon(A)+2$.
Proof.
$(1)\implies(2)$ Let $S$ be a leaf in $R(A)$. Since $\mathrm{dom}.\dim A\geq 2$, we know from Lemma 4.4 that $\tau^{-1}S$ is black. Then $P(S)$ is injective by Lemma 2.1. If $S$ is red, there is an exact sequence
$0\to P(\tau S)\to P(S)\to I(S)\to I(\tau^{-1}S)\to 0.$
Since $\mathrm{dom}.\dim A\geq 3$, the module $I(\tau^{-1}S)$ is projective and the sequence splits. This is impossible. Then $S$ is black.
$(2)\iff(3)$ This follows from Corollary 2.2.
$(3)\iff(4)$ Since $R(A)^{\prime}$ and $R(\varepsilon{A})$ are isomorphic quivers, the defect of $\varepsilon(A)$ is equal to the number of leaf vertices in $R(A)^{\prime}$. Then $\operatorname{def}A$ is equal to $\operatorname{def}\varepsilon(A)$ if and only if $S$ is the unique predecessor of its successor for every leaf $S$ in $R(A)$.
$(4)\implies(5)$ Let $T$ be a leaf in $R(A)$ and $S$ be the successor of $T$. Since $\tau^{-1}T$ is a black leaf, by (3.3) there is a non-split exact sequence
$0\to J(\tau^{-1}S)\to P(T)\to P(I(\tau^{-1}T))\to I(\tau^{-1}T)\to 0.$
Since $A$ satisfies (B), by Proposition 3.9 the projective injective objects in $\mathcal{F}$ and the projective injective $A$-modules contained in $\mathcal{F}$ coincide. Then
$\mathrm{codom}.\dim_{A}I(\tau^{-1}T)=\mathrm{codom}.\dim_{\mathcal{F}}J(\tau^{-1}S)+2.$
By Lemma 3.4 we have $\mathrm{dom}.\dim A=\mathrm{dom}.\dim\varepsilon(A)+2$.
$(5)\implies(1)$ Since $\mathrm{dom}.\dim\varepsilon(A)\geq 1$, we have $\mathrm{dom}.\dim A\geq 3$. ∎
Note that dominant dimension at least $3$ implies (B) and (B) for any Nakayama algebra. We have the following.
Corollary 4.8.
Let $A$ be a Nakayama algebra of dominant dimension at least $3$ and let $\Delta(S)$ be a vertex in $R(\varepsilon{A})$. Then
  1. (1)
    $\Delta(S)$ is black if and only if $S$ is black;
  2. (2)
    $\tau^{-1}_{\varepsilon}\Delta(S)$ is black if and only if $\tau^{-1}S$ is black.
Corollary 4.9.
Let $A$ be a Nakayama algebra of dominant dimension $d\geq 3$. Then $A$ is minimal Auslander-Gorenstein of selfinjective dimension $d$ if and only if $\varepsilon(A)$ is minimal Auslander-Gorenstein of selfinjective dimension d2d-2.
Proof.
This follows from Corollary 3.6 and Lemma 4.7. Compare [15, 4.4.2]. ∎
Now we restate and prove Theorem A. It gives a criterion to decide whether a Nakayama algebra is minimal Auslander-Gorenstein of odd selfinjective dimension.
Theorem 4.10.
Let $A$ be a Nakayama algebra and $R(A)$ be its resolution quiver. For any $m\geq 1$, the algebra $A$ is minimal Auslander-Gorenstein of selfinjective dimension 2m12m-1 if and only if
  1. (1)
    $\operatorname{dist}S=\operatorname{dist}A=m$ for every leaf $S$ in $R(A)$;
  2. (2)
    $|\gamma^{-1}(S)|\leq 1$ for every noncyclic vertex $S$ in $R(A)$;
  3. (3)
    $R(A)$ is connected of weight $1$;
  4. (4)
    $\tau^{-1}$ of all cyclic vertices in $R(A)$ are black.
Proof.
$\implies$" We prove by induction on $m$. If $m=1$, this is proven in Proposition 4.1. If $m\geq 2$, by Corollary 4.9 the algebra $\varepsilon(A)$ is minimal Auslander-Gorenstein of selfinjective dimension 2m32m-3. By induction hypothesis, the following statements hold for $\varepsilon(A)$.
  1. (1a)
    $\operatorname{dist}\Delta=\operatorname{dist}\varepsilon(A)=m-1$ for every leaf $\Delta$ in $R(\varepsilon A)$;
  2. (2a)
    $|\gamma_{\varepsilon}^{-1}(\Delta)|\leq 1$ for every noncyclic vertex $\Delta$ in $R(\varepsilon A)$;
  3. (3a)
    $R(\varepsilon A)$ is connected of weight $1$;
  4. (4a)
    $\tau^{-1}_{\varepsilon}$ of all cyclic vertices in $R(\varepsilon A)$ are black.
There is an injective map of resolution quivers from $R(\varepsilon{A})$ to $R(A)$ which preserves the weight. Then (1), (2) and (3) hold. By Corollary 4.8 we get (4).
[…]
There is an injective map of resolution quivers $\varphi$ from $R(\varepsilon{A})$ to $R(A)$ which preserves the weight. Then (1a), (2a) and (3a) hold. By Corollary 4.8 we obtain (4a).
By the induction hypothesis, the algebra $\varepsilon(A)$ is minimal Auslander-Gorenstein of selfinjective dimension 2m32m-3. Since the dominant dimension of $A$ is at least $3$, by Corollary 4.9 the algebra $A$ is minimal Auslander-Gorenstein of selfinjective dimension 2m12m-1. ∎
We restate Theorem B as follows. The proof is similar, we omit it here.
Theorem 4.11.
Let $A$ be a Nakayama algebra and $R(A)$ be its resolution quiver. For any $m\geq 1$, the algebra $A$ is minimal Auslander-Gorenstein of selfinjective dimension $2m$ if and only if
  1. (1)
    $\operatorname{dist}S=\operatorname{dist}A=m$ for every leaf $S$ in $R(A)$;
  2. (2)
    $|\gamma^{-1}(S)|\leq 1$ for every noncyclic vertex $S$ in $R(A)$;
  3. (3)
    $|\gamma^{-1}(S)|\leq 2$ for every cyclic vertex $S$ in $R(A)$;
  4. (4)
    all cyclic vertices in $R(A)$ are black.
For a Nakayama algebra $A$, denote by $\operatorname{cyc}A$ be the number of cyclic vertices in its resolution quiver. We have the following.
Corollary 4.12.
Let $A$ be a Nakayama algebra. Then $\operatorname{def}A\leq\operatorname{cyc}A$ provided that $A$ is minimal Auslander-Gorenstein of even selfinjective dimension.
5. Applications and examples
Let $A$ be a Nakayama algebra of rank $n$ with Kupisch series
$(c_{1},c_{2},\cdots,c_{n}).$
Denote by $\bar{A}$ the Nakayama algebra with Kupisch series
$(c_{1}+n,c_{2}+n,\cdots,c_{n}+n).$
It is direct to show that there is a color preserving isomorphism between $R(\bar{A})$ and $R(A)$. However, the weight of $R(\bar{A})$ is grater than the weight of $R(A)$.
By Theorem 4.11, we have the following.
Proposition 5.1.
Let $A$ be a Nakayama algebra and let $d$ be an even number. Then $A$ is minimal Auslander-Gorenstein of selfinjective dimension $d$ if and only if so is $\bar{A}$.
For any integer $n$, denote by $\mathbb{A}_{n}$ be the hereditary linear Nakayama algebra of rank $n$. For any $m\leq n$, denote by $N_{n}(m)$ be the quotient algebra of $\mathbb{A}_{n}$ via the ideal generated by all paths of length $m$. The Kupisch series of $N_{n}(m)$ is
$((m)^{n-m+1},m-1,\cdots,2,1).$
The algebra $N_{n}(m)$ has finite global dimension. If $m=2$, then $N_{n}(m)$ is higher Auslander of global dimension n1n-1.
Suppose $m>2$ and $N_{n}(m)$ is a higher Auslander algebra. Write
$n=qm+r$
such that $q,r$ are integers and $0\leq r\leq m-1$. Since $S(1)$ is cyclic and $\gamma^{-1}S(1)$ consists of $m$ simple $A$-modules, the global dimension of $N(n,r)$ is odd by Theorem 4.11. Since S(nr+1)S(n-r+1) is cyclic, then S(nr)S(n-r) is black by Theorem 4.10. We infer that $r=0$ and $m$ divides $n$.
If $m$ divides $n$, it is routine to check that $N_{n}(m)$ is higher Auslander of global dimension 2n/m12n/m-1.
Proposition 5.2.
The algebra $N_{n}(m)$ is higher Auslander if and only if $m=2$ or $m$ divides $n$. In both cases, the global dimension of $N_{n}(m)$ is 2n/m12n/m-1.
Following [15], for any Nakayama algebra $A$, there is a unique Nakayama algebra $\varepsilon^{-1}{A}$ of dominant dimension at least $3$ such that $\varepsilon(\varepsilon^{-1}{A})$ is isomorphic to $A$. Let $A$ be a Nakayama algebra of rank $n$ with Kupisch series
$c(A)=(c_{1},c_{2},\cdots,c_{n}).$
Denote by $d_{a}$ the maximum of $c_{a}-c_{a-1}$ and 0 for each $a\leq n$, and denote by $d$ the sum of all $d_{a}$. Let $\tilde{c}_{a}$ be a series of length $1+d_{a+1}$ such that
$(\tilde{c}_{a})_{\ell}=c_{a}+\sum_{i=1}^{c_{a}+\ell-1}d_{a+i}$
for each $\ell\leq 1+d_{a+1}$. Then $\varepsilon^{-1}{A}$ is a Nakayama algebra of rank $n+d$ with Kupisch series
$c(\varepsilon^{-1}A)=(\tilde{c}_{1},\tilde{c}_{2},\cdots,\tilde{c}_{n}).$
One sees that $\chi(\varepsilon^{-1}{A})$ is equal to the number of simple components of $A$.
Recall the following result in [15].
Proposition 5.3.
Let $d\geq 3$ and $\Lambda$ be a cyclic Nakayama algebra. Then $\Lambda$ is higher Auslander of global dimension $d$ if and only if $\Lambda$ is isomorphic to
$\varepsilon^{-1}(A)$
where $A$ is a higher Auslander cyclic Nakayama algebra of global dimension d2d-2 or
$\varepsilon^{-1}(A_{1}\oplus A_{2}\oplus\cdots\oplus A_{r})$
where $A_{i}$ is a higher Auslander linear Nakayama algebra of global dimension d2d-2 for $i\leq r$.
Similarly, we have the following.
Proposition 5.4.
Let $d\geq 3$ and $\Lambda$ be a linear Nakayama algebra. Then $\Lambda$ is higher Auslander of global dimension $d$ if and only if $\Lambda$ is isomorphic to
$\varepsilon^{-1}(\mathbb{A}_{1}\oplus A_{1}\oplus\cdots\oplus A_{r})$
where $A_{i}$ is a higher Auslander linear Nakayama algebra of global dimension d2d-2 for $i\leq r$.
Connected Nakayama algebras of which are higher Auslander of global dimension $1$ are hereditary. Connected Nakayama algebras of which are higher Auslander of global dimension $2$ are just the Auslander algebras of radical square zero connected Nakayama algebras. We can repeatedly apply Proposition 5.4 and 5.3 to give all connected Nakayama algebras of which are higher Auslander.
Example 5.5.
Let $A$ be a Nakayama algebra with Kupisch series
$(1,3,2,1,2,1).$
The Kupisch series of $\varepsilon^{-1}(A)$ is
$(3,3,3,4,3,2,2,2,1).$
Then $\varepsilon^{-1}(A)$ is higher Auslander of global dimension $3$.

### v2

@@
 Let $A$ be a Nakayama algebra.
-Using Ringel’s resolution quiver, we give a criterion to decide whether $A$ is minimal Auslander-Gorenstein.
+Using Ringel’s resolution quiver, we give a criterion to determine whether $A$ is a minimal Auslander-Gorenstein algebra.
 The criterion strongly relies on the parity of the selfinjective dimension of $A$.
@@
 1. Introduction
-Following Iyama [4], an algebra is said to be higher Auslander if its global dimension is finite and bounded by its dominant dimension.
-Higher Auslander algebras arise in cluster-tilting theory.
-Following Iyama and Solberg [5], an algebra is said to be minimal Auslander-Gorenstein provided that its global dimension is finite and bounded by its dominant dimension.
-Minimal Auslander-Gorenstein algebras are a special class of Auslander-Gorenstein algebras.
-Nakayama algebras are one of most basic classes of algebras in representation theory of algebras.
+Following Iyama [4], an algebra is said to be higher Auslander provided that its global dimension is finite and bounded by its dominant dimension.
+Higher Auslander algebras generalize the classical Auslander algebras.
+Following Iyama and Solberg [5], an algebra is said to be minimal Auslander-Gorenstein provided that its selfinjective dimension is finite and bounded by its dominant dimension.
+Minimal Auslander-Gorenstein algebras are the Gorenstein analogues of higher Auslander algebras.
+They are a special class of Auslander-Gorenstein algebras.
+Nakayama algebras are one of the most fundamental classes of algebras in the representation theory of algebras.
 An algebra is called a Nakayama algebra if it is a homomorphic image of the path algebra of an oriented cycle.
-A connected Nakayama algebras is either linear or else cyclic.
+The indecomposable modules over Nakayama algebras are uniserial.
+A connected Nakayama algebra is either linear or cyclic.
 The resolution quiver for Nakayama algebras is introduced by Ringel [10].
@@
 A resolution quiver criterion to detect the finiteness of global dimension and the finiteness of selfinjective dimension for Nakayama algebras is given in [17].
-The syzygy filtered algebras for cyclic Nakayama algebras is introduced by Sen [12].
-It plays a role in constructing cyclic Nakayama algebras which are higher Auslander in [13].
-The reversed syzygy filtered algebras for all Nakayama algebras is introduced in [15].
-It provides a systematic construction of cyclic Nakayama algebras which are minimal Auslander-Gorenstein and higher Auslander.
-Inspired by their work, we study syzygy filtered algebras for all Nakayama algebras.
-Then we give a resolution quiver criterion to decide whether a Nakayama algebra is minimal Auslander-Gorenstein.
+The syzygy filtered algebra for cyclic Nakayama algebras is introduced by Sen [13].
+It plays a role in constructing cyclic Nakayama algebras which are higher Auslander algebras [14].
+The reversed syzygy filtered algebra for all Nakayama algebras is introduced in [16].
+It provides a systematic method to construct cyclic Nakayama algebras which are higher Auslander and minimal Auslander-Gorenstein.
+A classification of defect one cyclic Nakayama algebras is given in [7], and a classification of concave linear Nakayama algebras which are higher Auslander is given in [12].
+Inspired by their work, we study the syzygy filtered algebra for all Nakayama algebras.
+Then we give a resolution quiver criterion to determine whether a Nakayama algebra is a minimal Auslander-Gorenstein algebra.
 This criterion strongly depends on the parity of the selfinjective dimension.
-Let $A$ be a Nakayama algebra with translation $\tau$.
-For an $A$-module $M$, denote by $P(M)$ the projective cover and by $I(M)$ the injective envelope of $M$.
+Let $A$ be a Nakayama algebra with translation $\tau$ on the simple $A$-modules.
+For an $A$-module $M$, denote by $\ell(M)$ the composition length, by $P(M)$ the projective cover, and by $I(M)$ the injective envelope of $M$.
 The resolution quiver $R(A)$ of $A$ is defined as follows.
-The vertex set of $R(A)$ is the isoclasses of all simple $A$-modules, and there is a unique arrow from $S$ to $\tau\operatorname{soc}P(S)$ for each simple $A$-module $S$.
-A simple $A$-module $S$ is said to be black if the projective dimension of $S$ is not equal to $1$, and $S$ is said to be red otherwise.
-This assigns colors to the vertices of resolution quivers.
-The weight of a cycle $C$ in $R(A)$ is the sum of $c_{a}/n$, where $a$ runs through all vertices in $C$ and $n$ is rank of $A$.
-The weight is an integer and all cycles in $R(A)$ have the same weight.
-The weight of any cycle in $R(A)$ is called the weight of $R(A)$.
-A simple $A$-module $S$ said to be cyclic if $S$ lies on cycle in $R(A)$.
-For any simple $A$-module $S$, denote by $\operatorname{dist}S$ the minimal integer $d$ such that the $d$-th successor of $S$ is cyclic in $R(A)$.
-Let $\operatorname{dist}A$ be the maximum of the numbers $\operatorname{dist}S$, where $S$ runs through all simple $A$-modules.
+The vertices of $R(A)$ are the isoclasses of all simple $A$-modules, and there is a unique arrow from $S$ to $\tau\operatorname{soc}P(S)$ for each simple $A$-module $S$.
+A simple $A$-module $S$ is said to be red if the projective dimension of $S$ is equal to $1$, and black otherwise.
+This assigns a color to each vertex of the resolution quiver.
+A simple $A$-module $S$ is said to be cyclic if $S$ lies on a cycle in $R(A)$, and a leaf if there is no arrow in $R(A)$ ending at $S$.
+For any simple $A$-module $S$, we denote by $\operatorname{dist}S$ the minimal integer $d$ such that the $d$-th successor of $S$ is cyclic.
+Let $\operatorname{dist}A$ be the maximum of $\operatorname{dist}S$ over all simple $A$-modules $S$.
+The weight of $R(A)$ is the sum of $\ell(P(S))/nc$, where $n$ is the rank of $A$, $c$ is the number of cycles in $R(A)$, and $S$ runs through all cyclic vertices in $R(A)$.
+It turns out that a Nakayama algebra has finite global dimension if and only if its resolution quiver is connected of weight $1$.
 The following is the main result of this paper.
 Theorem A.
-Let $A$ be a Nakayama algebra and let $R(A)$ be the resolution quiver of $A$. Then $A$ is minimal Auslander-Gorenstein of odd selfinjective dimension if and only if
+Let $A$ be a Nakayama algebra and $R(A)$ its resolution quiver. Then $A$ is minimal Auslander-Gorenstein of odd selfinjective dimension if and only if
   1. (1)
-    $\operatorname{dist}S=\operatorname{dist}A$ for every leaf $S$ in $R(A)$;
+    $\operatorname{dist}S=\operatorname{dist}A>0$ for every leaf $S$ in $R(A)$;
   2. (2)
@@
 Theorem B.
-Let $A$ be a Nakayama algebra and let $R(A)$ be the resolution quiver of $A$. Then $A$ is minimal Auslander-Gorenstein of even selfinjective dimension if and only if
+Let $A$ be a Nakayama algebra and $R(A)$ its resolution quiver. Then $A$ is minimal Auslander-Gorenstein of even selfinjective dimension if and only if
   1. (1)
-    $\operatorname{dist}S=\operatorname{dist}A$ for every leaf $S$ in $R(A)$;
+    $\operatorname{dist}S=\operatorname{dist}A>0$ for every leaf $S$ in $R(A)$;
   2. (2)
@@
   3. (3)
-    every cyclic vertex has at most $2$ predecessor in $R(A)$;
+    every cyclic vertex has at most $2$ predecessors in $R(A)$;
   4. (4)
@@
 All modules are finite dimensional left modules unless otherwise stated.
+We refer to [1] for basics on finite dimensional algebras and their representations.
 The paper is organized as follows.
-In Section 2, we recall some basic properties of Nakayama algebras.
-In Section 3, we study syzygy filtered algebra for all Nakayama algebras.
-Then we prove the criterion in Section 4.
-Some applications and examples is given in Section 5.
+In Section 2, we recall some basic properties of Nakayama algebras and their resolution quivers.
+In Section 3, we investigate the syzygy filtered algebra for all Nakayama algebras.
+Then we prove the main result in Section 4.
+Some applications and examples are given in Section 5.
 Acknowledgements
-The author is grateful to Professor Shijie Zhu for his comments and suggestions. This work is supported by the National Natural Science Foundation of China (No. 1257011205).
+The author is grateful to Professor Shijie Zhu for his comments and suggestions. This work is supported by the National Natural Science Foundation of China (No. 12571036 and 11801141).
 2. Preliminary
-We refer to [1] for quivers and their representations.
 Given an integer $n$, we denote by $\Delta_{n}$ the quiver on $n$ vertices containing a single directed cycle through all vertices; see Figure 1.
-Denote by $k\Delta_{n}$ the path algebra of the quiver $\Delta_{n}$.
+Denote by $k\Delta_{n}$ the path algebra of $\Delta_{n}$.
 112233nn
@@
 Let $A$ be a Nakayama algebra of rank $n$.
-For each vertex $a$ in $\Delta_{n}$, denote by $S(a)$ the simple $A$-module, $P(a)$ the projective $A$-module and $I(a)$ the injective $A$-module at $a$.
-Set $\tau S(a)=S(b)$ if there is an arrow from $a$ to $b$ in $\Delta_{n}$.
-The map $\tau$ yields a translation of simple $A$-modules.
-If $A$ is a cyclic Nakayama algebra, then $\tau$ coincides with the Auslander-Reiten translation.
+For each vertex $a$ in $\Delta_{n}$, denote by $S(a)$ the simple $A$-module, by $P(a)$ the projective $A$-module, and by $I(a)$ the injective $A$-module at $a$.
+Set $\tau(a)=b$ if there is an arrow from $a$ to $b$ in $\Delta_{n}$.
+This induces a translation $\tau$ on the simple $A$-modules.
+If $A$ is a cyclic Nakayama algebra, then $\tau$ coincides with the Auslander-Reiten translation [1].
 For each vertex $a$ in $\Delta_{n}$, denote by $c_{a}$ the composition length of $P(a)$. The series
-$c(A)=(c_{a},c_{\tau a},\cdots,c_{\tau^{n-1}a})$
-is called a Kupisch series of $A$. Two Nakayama algebras $A$ and $B$ are isomorphic if and only if $c(A)$ and $c(B)$ are equal up to a power of cyclic permutation.
-Denote by $R(A)$ the resolution quiver of $A$. The vertices of $R(A)$ are the isoclasses of all simple $A$-modules, and there is a unique arrow from $S$ to
+$c(A)=(c_{1},c_{2},\cdots,c_{n})$
+is called a Kupisch series of $A$. Two Nakayama algebras are isomorphic if and only if their Kupisch series are equal up to cyclic permutation.
+Following [10], the resolution quiver $R(A)$ of $A$ is the successor graph of the function
 $\gamma(S)=\tau\operatorname{soc}P(S)$
-for each simple $A$-module $S$; see also [2].
-A simple $A$-module $S$ is said to be black if the projective dimension of $S$ is not equal to $1$, and $S$ is said to be red otherwise.
-This assigns colors to the vertices of $R(A)$.
+on the set of all isoclasses of simple $A$-modules; see also [2].
+A simple $A$-module $S$ is said to be red if the projective dimension of $S$ is equal to $1$, and black otherwise.
+A projective $A$-module $P$ is said to be minimal if its top module is a direct sum of black simple $A$-modules.
 The following is well known.
@@
   2. (2)
-    $\gamma(\tau S)=\gamma(S)$, provided that $S$ is red.
+    $\ell(P(S))=\ell(P(\tau S))+1$ if and only if $S$ is red.
 Corollary 2.2.
 Let $S$ be a simple $A$-module. Then $S$ and $\tau^{-1}S$ are black if and only if $\gamma^{-1}(\gamma S)=\{S\}$.
-We need the following; see [16, 1.1] and [11, 3.10].
+We need the following; see [17, 1.1] and [11, 3.10].
 Lemma 2.3.
@@
     $2\operatorname{dist}A-1\leq\mathrm{gl}.\dim A\leq 2\operatorname{dist}A$, provided that $\mathrm{gl}.\dim A$ is finite.
-An algebra $A$ is said to be Gorenstein if the left selfinjective dimension and right selfinjective dimension of $A$ are finite.
-For any Gorenstein algebra, two selfinjective dimensions coincide; see [17].
+An algebra $A$ is said to be Gorenstein if its left selfinjective dimension and right selfinjective dimension are finite.
+If $A$ is a Gorenstein algebra, then its left and right selfinjective dimensions are equal; see [18].
 It is known that a Nakayama algebra is Gorenstein if and only if its left selfinjective dimension is finite.
-We need the following; see [16, 1.2] and [12, 4.13].
+We need the following; see [17, 1.2] and [13, 4.13].
 Lemma 2.4.
@@
   1. (1)
-    AA is Gorenstein if and only if all cycles in $R(A)$ are black;
+    AA is Gorenstein if and only if all cyclic vertices in $R(A)$ are black;
   2. (2)
     $\mathrm{id}_{A}\,A=2\operatorname{dist}A$, provided that $\mathrm{id}_{A}\,A$ is finite.
-Let $A$ be a Nakayama algebra and $R(A)$ be its resolution quiver. A vertex $S$ in $R(A)$ is said to be a leaf if there is no arrow ending with $S$.
-We need the following; see [6, 3.1].
+Let $A$ be a Nakayama algebra. A simple $A$-module $S$ is said to be a leaf if there is no arrow in $R(A)$ ending at $S$. We need the following; see [6, 3.1].
 Lemma 2.5.
-Let $S$ be a simple $A$-module. Then $\mathrm{id}_{A}\,S=1$ if and only if $S$ is a leaf in $R(A)$.
+Let $S$ be a simple $A$-module. Then
+  1. (1)
+    $I(S)$ is projective if and only if $\tau S$ is a non-leaf in $R(A)$;
+  2. (2)
+    $\mathrm{id}_{A}\,S=1$ if and only if $S$ is a leaf in $R(A)$.
 3. Syzygy filtered algebra
-In the present section, we study the syzygy filtered algebra for all Nakayama algebras and give some necessary results for the proof of the main result.
-Let $A$ be a Nakayama algebra of rank $n$ and let $R(A)$ be the resolution quiver of $A$.
-Following [11, A.6] the map $\gamma$ induces a one-to-one correspondence between the set of all black vertices and the set of all non-leaf vertices in $R(A)$.
+In this section, we study the syzygy filtered algebra for all Nakayama algebras. This generalizes Sen’s construction for cyclic Nakayama algebras.
+Let $A$ be a Nakayama algebra and $R(A)$ its resolution quiver.
+By [11, A.6] the map $\gamma$ induces a one-to-one correspondence between the set of all black vertices and the set of all non-leaf vertices in $R(A)$.
 For any non-leaf vertex $S$ in $R(A)$, let $T$ be the unique black vertex such that $\gamma(T)=S$.
 Consider the exact sequence
-$0\to\Delta(S)\to P(\tau T)\to P(T)\to T\to 0.$ (3.1)
-The base module $\Delta(S)$ is the quotient module of $P(S)$ of minimal length such that the translation of its socle is a non-leaf vertex.
+$0\to\Delta(S)\to P(\tau T)\overset{\alpha}{\to}P(T)\to T\to 0,$ (3.1)
+where $\alpha\in A$ is the arrow starting at $T$.
+The kernel $\Delta(S)$ is called a base module.
+It is the quotient module of $P(S)$ of minimal length such that the translation of its socle is a non-leaf vertex.
 For any module $M$, denote by $\Omega(M)$ the syzygy module of $M$. Then we have
 $\Delta(S)=\begin{cases}P(S),&\text{if}\;\mathrm{pd}_{A}\,T\in\{0,2\},\\ \Omega^{2}(T),&\text{otherwise}.\end{cases}$ (3.2)
-Following [12, 14], the base set
+Following [13, 15, 11], the base set
 $\mathcal{B}=\{\Delta(S)\mid\text{$S$ is a non-leaf vertex in $R(A)$}\}$
@@
 The filtration category $\mathcal{F}$ of $\mathcal{B}$ is an exact Abelian subcategory of $\operatorname{mod}A$.
-The category $\mathcal{F}$ contains the projective cover $\mathcal{P}$ of all base modules and the second syzygy of all finite dimensional $A$-modules.
-The syzygy filtered algebra $\varepsilon(A)$ is the endomorphism algebra of $\mathcal{P}$, that is,
+The category $\mathcal{F}$ contains the projective cover of every base module and the second syzygy of every finite dimensional $A$-module.
+Denote by $\mathcal{P}$ the projective covers of all base modules. The syzygy filtered algebra $\varepsilon(A)$ of $A$ is the endomorphism algebra of $\mathcal{P}$, that is,
 $\varepsilon(A)=\operatorname{End}_{A}(\bigoplus\nolimits_{P\in\mathcal{P}}P).$
-It is a Nakayama algebra of rank at most $n$.
+It is a Nakayama algebra of rank less than or equal to the rank of $A$.
 The equality holds if and only if $A$ is selfinjective.
-The category $\operatorname{mod}\varepsilon(A)$ is equivalent to $\mathcal{F}$.
-The algebra $\varepsilon(A)$ has finite global dimension if and only if $A$ has finite global dimension.
-For a Nakayama algebra $A$, denote by $\chi(A)$ the number of isoclasses of simple projective $A$-modules. By (3.2), we have the following.
+The category $\operatorname{mod}\varepsilon(A)$ of all finite dimensional $\varepsilon(A)$-modules is equivalent to $\mathcal{F}$.
+The syzygy filtered algebra $\varepsilon(A)$ has finite global dimension if and only if $A$ has finite global dimension.
+Given a Nakayama algebra $A$, denote by $\chi(A)$ the number of isoclasses of simple projective $A$-modules.
+Note that $\chi(A)$ equals the number of occurrences of $1$ in the Kupisch series of $A$.
+For any integer $i\geq 0$, denote by $\varepsilon^{i}(A)$ the $i$-th iterated syzygy filtered algebra of $A$.
+By (3.2), we have the following.
 Lemma 3.1.
-Let $A$ be a Nakayama algebra. Then $\chi(\varepsilon{A})-\chi(A)$ is equal to the number of isoclasses of simple $A$-modules of projective dimension $2$.
-For any integer $i\geq 0$, denote by $\varepsilon^{i}(A)$ the $i$-th iterated syzygy filtered algebra of $A$. We have the following.
-Corollary 3.2.
-For any $i\geq 1$, $\chi(\varepsilon^{i}{A})-\chi(\varepsilon^{i-1}A)$ is equal to the number of isoclasses of simple $A$-modules of projective dimension $2i$.
-There is an integer $r$ such that $\varepsilon^{r}(A)$ is a selfinjective Nakayama algebra.
-Then $\chi(\varepsilon^{r}{A})$ is equal to the number of isoclasses of simple $A$-modules of even projective dimension.
-This gives an alternative proof of Madsen’s criterion [6], which states that a Nakayama algebra has finite global dimension if and only if it has a simple module of even projective dimension.
-The projective objects in $\mathcal{F}$ are precisely the projective $A$-modules which are contained in $\mathcal{F}$. However, the injective objects in $\mathcal{F}$ are more complicated.
+Let $A$ be a Nakayama algebra and $i\geq 1$. Then $\chi(\varepsilon^{i}{A})-\chi(\varepsilon^{i-1}A)$ is equal to the number of isoclasses of simple $A$-modules of projective dimension $2i$.
+For any $i\geq 1$, the rank of $\varepsilon^{i}(A)$ is no more than the rank of $\varepsilon^{i-1}(A)$.
+There is an integer $m$ such that $\varepsilon^{m}(A)$ is selfinjective.
+Then $\chi(\varepsilon^{m}{A})$ is equal to the number of isoclasses of simple $A$-modules of even projective dimension.
+This gives an alternative proof of Madsen’s criterion [6], which states that a Nakayama algebra $A$ has finite global dimension if and only if there is a simple $A$-module of even projective dimension.
 Following [7], the defect $\operatorname{def}A$ of an algebra $A$ is the number of isoclasses of indecomposable injective non-projective $A$-modules.
-If $A$ is a Nakayama algebra, then $\operatorname{def}A$ is equal to the number of red vertices in its resolution quiver.
+If $A$ is a Nakayama algebra, the defect of $A$ is equal to the number of red vertices in the resolution quiver of $A$.
 By (3.2) we have following.
 Lemma 3.2.
-For any $i\geq 0$, $\operatorname{def}\varepsilon^{i}(A)$ is equal to the number of isoclasses of simple $A$-modules of projective dimension $2i+1$.
-Let $S$ be a non-leaf vertex in $R(A)$. By [11, A.6] there is a unique vertex $T$ such that $P(T)$ is injective and $\gamma(T)=S$. Consider the exact sequence
-$0\to J(\tau^{-1}S)\to P(T)\to P(I(\tau^{-1}T))\overset{f}{\to}I(\tau^{-1}T)\to 0.$ (3.3)
-We have the following; compare [12, 3.18].
+Let $A$ be a Nakayama algebra and $i\geq 0$. Then $\operatorname{def}\varepsilon^{i}(A)$ is equal to the number of isoclasses of simple $A$-modules of projective dimension $2i+1$.
+The projective objects in $\mathcal{F}$ are precisely the projective $A$-modules contained in $\mathcal{F}$. The injective objects in $\mathcal{F}$ are more complicated. Let us give a description.
+Let $S$ be a non-leaf vertex in $R(A)$. By [11, A.6] there is a unique vertex $T$ such that $\tau^{-1}T$ is black and $\gamma(T)=S$. Consider the exact sequence
+$0\to J(\tau^{-1}S)\to P(T)\overset{w}{\to}P(I(\tau^{-1}T))\to I(\tau^{-1}T)\to 0,$ (3.3)
+where $w\in A$ is the concatenation of the maximal nonzero path ending at $\tau^{-1}T$ and the arrow starting at $\tau^{-1}T$.
+We have the following; compare [13, 3.18].
 Lemma 3.3.
@@
 Proof.
-If $T$ is a non-leaf vertex, the module $I(\tau^{-1}T)$ is projective. The sequence (3.3) splits and $J(\tau^{-1}S)$ is an indecomposable injective $A$-module contained in $\mathcal{F}$.
-If $T$ is a leaf, then $I(\tau^{-1}T)$ is not projective and $\operatorname{Ker}f$ is nonzero.
-Since $P(T)$ is an injective $A$-module, $J(\tau^{-1}S)$ is nonzero.
-Then $J(\tau^{-1}S)$ is an indecomposable object in $\mathcal{F}$.
-Since $\operatorname{Ker}f$ is a proper submodule of a base module and $P(T)$ is an injective $A$-module, we know that $J(\tau^{-1}S)$ is an injective object in $\mathcal{F}$.
-The number of isoclasses of injective objects in $\mathcal{F}$ is equal to the number of non-leaf vertices in $R(A)$.
-If $S$ and $S^{\prime}$ are not isomorphic, $J(\tau^{-1}S)$ and $J(\tau^{-1}S^{\prime})$ are not isomorphic.
-Then the first conclusion holds.
-Note that $J(\tau^{-1}S)$ is a non-projective object in $\mathcal{F}$ if and only if $\Delta(S)$ is a leaf in $R(\varepsilon{A})$. This happens if and only if the injective dimension of $S$ is $3$. ∎
-The translation $\tau$ induces a translation $\tau_{\varepsilon}$ of simple objects in $\mathcal{F}$. For any vertex $\Delta(S)$ in $R(\varepsilon{A})$, by [11, B.4] we have
+If $T$ is a non-leaf vertex, then $I(\tau^{-1}T)$ is projective and $w$ is zero. It follows that $J(\tau^{-1}S)$ is an indecomposable injective $A$-module contained in $\mathcal{F}$.
+If $T$ is a leaf, then $I(\tau^{-1}T)$ is not projective and $w$ is nonzero.
+Since $P(T)$ is an injective $A$-module, $J(\tau^{-1}S)$ is an indecomposable object in $\mathcal{F}$.
+Since the image of $w$ is a proper submodule of a base module, we infer that $J(\tau^{-1}S)$ is an injective object in $\mathcal{F}$.
+If $S$ and $S^{\prime}$ are not isomorphic, the socle modules of $J(\tau^{-1}S)$ and $J(\tau^{-1}S^{\prime})$ are not isomorphic.
+Then $J(\tau^{-1}S)$ and $J(\tau^{-1}S^{\prime})$ are not isomorphic.
+Note that the number of isoclasses of indecomposable injective objects in $\mathcal{F}$ is equal to the number of non-leaf vertices in $R(A)$.
+Then the assignment is a bijection.
+By Lemma 2.5 $J(\tau^{-1}S)$ is a non-projective object in $\mathcal{F}$ if and only if $\Delta(S)$ is a leaf in $R(\varepsilon{A})$.
+This happens if and only if the injective dimension of $S$ is $3$.
+∎
+The translation $\tau$ induces a translation $\tau_{\varepsilon}$ one the simple objects in $\mathcal{F}$. For any vertex $\Delta(S)$ in $R(\varepsilon{A})$, by [11, B.4] we have
 $\gamma_{\varepsilon}(\Delta(S))=\tau_{\varepsilon}\operatorname{soc}_{\mathcal{F}}P(\Delta(S))=\Delta(\gamma S).$
-The assignment $\Delta(S)\mapsto S$ yields an injective map
-$\varphi\colon R(\varepsilon{A})\to R(A)$
-of resolution quivers. Denote by $R(A)^{\prime}$ the image of $\varphi$. We mention that $\varphi$ preserves the weight of any cycle. However, $\varphi$ may not preserve the vertex colors.
-A vertex $T$ in $R(A)$ is said to be *-black if $\tau^{-1}T$ is black, and $T$ is said to be *-red if $\tau^{-1}T$ is red.
-For each leaf $S$ in $R(A)^{\prime}$, there is a unique black leaf $T_{0}$ in $R(A)$ such that $\gamma(T_{0})=S$ and a unique *-black leaf $T_{1}$ in $R(A)$ such that $\gamma(T_{1})=S$.
-For any resolution quiver $R$, denote by $\mathrm{Leaf}\mbox{-}R,\mathrm{BLeaf}\mbox{-}R$ and $\mathrm{*BLeaf}\mbox{-}R$ the set of all leaf vertices, black leaf vertices and *-black leaf vertices, respectively.
+The assignment $\Delta(S)\mapsto S$ yields an injective map of resolution quivers
+$\varphi\colon R(\varepsilon{A})\to R(A).$
+We denote by $R(A)^{\prime}$ the image of $\varphi$.
+Then $R(A)^{\prime}$ is obtained from $R(A)$ by removing all leaf vertices.
+Note that $\varphi$ preserves the weight.
+However, $\varphi$ may not preserve the vertex colors.
+A simple $A$-module $S$ is called *-black if $\tau^{-1}S$ is black, and *-red otherwise.
+For any resolution quiver $R$, denote by $\mathrm{Leaf}\mbox{-}R,\mathrm{BLeaf}\mbox{-}R$, and $\mathrm{*BLeaf}\mbox{-}R$ the set of all leaf vertices, black leaf vertices, and *-black leaf vertices, respectively.
+For each leaf $S$ in $R(A)^{\prime}$, there is a unique black leaf $T_{0}$ in $R(A)$ such that $\gamma(T_{0})=S$, and a unique *-black leaf $T_{1}$ in $R(A)$ such that $\gamma(T_{1})=S$.
 There is an injective map
-$\varphi_{0}\colon\mathrm{Leaf}\mbox{-}R(A)^{\prime}\to\mathrm{BLeaf}\mbox{-}R(A)$
-sending $S$ to $T_{0}$ and an injective map
-$\varphi_{1}\colon\mathrm{Leaf}\mbox{-}R(A)^{\prime}\to\mathrm{*BLeaf}\mbox{-}R(A)$
-sending $S$ to $T_{1}$. Consider the following two conditions.
+$\varphi_{0}\colon\mathrm{Leaf}\mbox{-}R(\varepsilon{A})\to\mathrm{BLeaf}\mbox{-}R(A)$
+sending $\Delta(S)$ to $T_{0}$ and an injective map
+$\varphi_{1}\colon\mathrm{Leaf}\mbox{-}R(\varepsilon{A})\to\mathrm{*BLeaf}\mbox{-}R(A)$
+sending $\Delta(S)$ to $T_{1}$. Consider the following two conditions.
   1. (B)
-    For each black leaf in $R(A)$, each predecessor of its successor is a leaf.
+    For each black leaf in $R(A)$, its successor is a leaf in $R(A)^{\prime}$.
   2. (B)
-    For each *-black leaf in $R(A)$, each predecessor of its successor is a leaf.
+    For each *-black leaf in $R(A)$, its successor is a leaf in $R(A)^{\prime}$.
 Then $A$ satisfies (B) if and only if the map $\varphi_{0}$ is bijective, and $A$ satisfies (B) if and only if the map $\varphi_{1}$ is bijective.
@@
   2. (2)
-    $\varphi$ preserves vertex colors;
+    $\varphi$ preserves the vertex colors;
   3. (3)
-    The number of red vertices in $R(\varepsilon{A})$ and that $R(A)^{\prime}$ are equal.
+    the number of red vertices in $R(\varepsilon{A})$ and $R(A)^{\prime}$ is equal;
+  4. (4)
+    any minimal projective object in $\mathcal{F}$ is a minimal projective $A$-module.
 Proof.
 $(1)\iff(3)$ Recall that the number of red vertices is equal to the number of leaf vertices in any resolution quiver. We have
-$|\mathrm{Red}\mbox{-}R(A)^{\prime}|=|\mathrm{Red}\mbox{-}R(A)|-|\mathrm{RLeaf}\mbox{-}R(A)|=|\mathrm{BLeaf}\mbox{-}R(A)|$
+$|\mathrm{Red}\mbox{-}R(\varepsilon{A})|=|\mathrm{Leaf}\mbox{-}R(\varepsilon{A})|\leq|\mathrm{BLeaf}\mbox{-}R(A)|$
 and
-$|\mathrm{Red}\mbox{-}R(\varepsilon{A})|=|\mathrm{Leaf}\mbox{-}R(\varepsilon{A})|=|\mathrm{Leaf}\mbox{-}R(A)^{\prime}|\leq|\mathrm{BLeaf}\mbox{-}R(A)|.$
+$|\mathrm{BLeaf}\mbox{-}R(A)|=|\mathrm{Leaf}\mbox{-}R(A)|-|\mathrm{RLeaf}\mbox{-}R(A)=|\mathrm{Red}\mbox{-}R(A)^{\prime}|.$
 The equality holds if and only if $\varphi_{0}$ is a bijection.
 $(2)\iff(3)$ Let $P$ be an indecomposable projective object in $\mathcal{F}$.
-If $P$ is minimal in $\operatorname{mod}A$, so is $P$ in $\mathcal{F}$.
+If $P$ is minimal projective in $\operatorname{mod}A$, then $P$ is minimal projective in $\mathcal{F}$.
 This shows that $\varphi$ preserves red vertices.
-Then $\varphi$ preserves vertex colors if and only if the numbers of red vertices in $R(\varepsilon{A})$ and $R(A)^{\prime}$ are equal.
-∎
+Then $\varphi$ preserves the vertex colors if and only if the number of red vertices in $R(\varepsilon{A})$ and $R(A)^{\prime}$ is equal.
+$(2)\iff(4)$ This is obvious. ∎
+Similarly, we have the following.
+Proposition 3.5.
+Let $A$ be a Nakayama algebra. The following are equivalent.
+  1. (1)
+    AA satisfies (B);
+  2. (2)
+    $\varphi$ preserves the vertex *-colors;
+  3. (3)
+    the number of *-red vertices in $R(\varepsilon{A})$ and $R(A)^{\prime}$ is equal;
+  4. (4)
+    any injective projective object in $\mathcal{F}$ is an injective projective $A$-module.
 Corollary 3.6.
-Let $A$ be a Nakayama algebra satisfying (B). Then $A$ is Gorenstein if and only if $\varepsilon(A)$ is Gorenstein. Moreover, we have
+Let $A$ be a Nakayama algebra satisfying (B). Then $A$ is Gorenstein if and only if $\varepsilon(A)$ is Gorenstein. Furthermore, we have
 $\mathrm{id}_{A}\,A=\mathrm{id}_{\varepsilon(A)}\,\varepsilon(A)+2$ (3.4)
@@
 Proof.
-Since $A$ satisfies (B), then $\varphi$ preserves vertex colors.
+Since $A$ satisfies (B), then $\varphi$ preserves the vertex colors.
 Recall that $A$ has finite global dimension if and only if $\varepsilon(A)$ has finite global dimension.
 Then by Lemma 2.4 the algebra $A$ is Gorenstein if and only if $\varepsilon(A)$ is Gorenstein.
-Note that the selfinjective dimension of $A$ is not $1$.
-The equation (3.4) follows from [12, 3.23].
-∎
+If $\mathrm{id}_{A}\,A=1$, then $A$ is hereditary by Lemma 2.4. There is a black leaf in $R(A)$. This is impossible. The equation (3.4) follows from [13, 3.23]. ∎
 Example 3.7.
@@
 The color of $1$ is not preserved.
-Similarly, we have the following.
-Proposition 3.9.
-Let $A$ be a Nakayama algebra. The following are equivalent.
-  1. (1)
-    AA satisfies (B);
-  2. (2)
-    $\varphi$ preserves vertex *-colors;
-  3. (3)
-    The number of *-red vertices in $R(\varepsilon{A})$ and $R(A)^{\prime}$ are equal.
 Example 3.9.
@@
 $0\to J(1)\to P(3)\to P(1)\to I(2)\to 0.$
-Note that $J(4)$ is isomorphic to $P(1)$ and $J(1)$ is an indecomposable injective non-projective object in $\mathcal{F}$.
+Note that $J(4)$ is an injective projective $A$-module, and $J(1)$ is a non-projective $A$-module.
 4. Proof of the main result
-In this section, we study the dominant dimensions of Nakayama algebras. Then we prove the main result of the present paper.
+In this section, we study the dominant dimension of Nakayama algebras by means of resolution quivers. Then we prove the main result of the present paper.
 Let $A$ be an algebra and $M$ be an $A$-module. Take a minimal injective resolution
@@
 The codominant dimension $\mathrm{codom}.\dim_{A}M$ of $M$ is the maximal integer $n$ or $\infty$ such that $P_{j}$ is injective for all $j<n$.
-The left dominant dimension of $A$ is equal to the right dominant dimension of $A$; see [8].
-It is known that any Nakayama algebra has dominant dimension at least $1$.
-For an integer $d\geq 1$, an algebra $A$ is called higher Auslander if
+The left and right dominant dimensions of $A$ are equal; see [8].
+It is known that the dominant dimension of any Nakayama algebra is at least $1$.
+Following [4, 5], for an integer $d\geq 1$, an algebra $A$ is called higher Auslander if
 $\mathrm{dom}.\dim A=d=\mathrm{gl}.\dim A,$
@@
 $\mathrm{dom}.\dim A=d=\mathrm{id}_{A}\,A.$
+Higher Auslander algebras are just the minimal Auslander-Gorenstein algebras of finite global dimension.
 For algebras of selfinjective dimension $1$, we have the following.
 Proposition 4.1.
-Let $A$ be a Nakayama algebra and $R(A)$ be its resolution quiver. Then $A$ is minimal Auslander-Gorenstein of selfinjective dimension $1$ if and only if
+Let $A$ be a Nakayama algebra and $R(A)$ its resolution quiver. Then the selfinjective dimension of $A$ is equal to $1$ if and only if
   1. (1)
@@
   3. (3)
-    $\tau^{-1}$ of all cyclic vertices in $R(A)$ are black.
+    all cyclic vertices in $R(A)$ are *-black.
 Proof.
-$\implies$" Since $\mathrm{id}_{A}\,A=1$, we have $\mathrm{gl}.\dim A=1$ by Lemma 2.4. Then by Lemma 2.3 $R(A)$ is connected of weight $1$. There is a direct sum decomposition
-$A=A_{1}\oplus A_{2}\oplus\cdots\oplus A_{r}$
-where $A_{i}$ is a hereditary linear Nakayama algebra for all $i\leq r$.
+$\implies$" Since $\mathrm{id}_{A}\,A=1$, by Lemma 2.4 we have $\mathrm{gl}.\dim A=1$. There is a direct sum decomposition of algebras
+$A=A_{1}\oplus A_{2}\oplus\cdots\oplus A_{r},$
+where $A_{i}$ is a hereditary linear Nakayama algebra for each $i\leq r$.
+By Lemma 2.3 $\operatorname{dist}A=1$ and $R(A)$ is connected of weight $1$.
 Let $S$ be a cyclic vertex in $R(A)$.
@@
 […]
-$0\to\Delta(S)\to P(\tau T)\to P(T)\to T\to 0$
-such that $\tau^{-1}S$ is black.
-Then $P(S)$ is injective by Lemma 2.1.
-Since $\mathrm{gl}.\dim A$ is at most $2$, by (3.2) we have $\Delta(S)=P(S)$.
-Then the sequence splits and $T$ is projective.
-We conclude that the selfinjective dimension of $A$ is $1$.
+$0\to\Delta(S)\to P(\tau T)\overset{\alpha}{\to}P(T)\to T\to 0.$
+Since $R(A)$ is connected of weight $1$, by Lemma 2.3 the global dimension of $A$ is at most $2$.
+It follows from (3.2) that $\Delta(S)=P(S)$.
+Then $\alpha$ is zero and $T$ is projective.
+We conclude that the selfinjective dimension of $A$ is equal to $1$.
 ∎
 Remark 4.2.
-Let $A$ be a Nakayama algebra of global dimension $1$. Since not all components of $A$ are simple, there is at least one red cyclic vertex in $R(A)$.
+Let $A$ be a Nakayama algebra of global dimension $1$. Since not all connected component of $A$ are simple, there exists at least one black leaf in $R(A)$.
 Example 4.3.
-Let $A$ be a Nakayama algebra with Kupisch series $(1,3,2,1)$.
-Then $A$ is the direct sum of two hereditary linear Nakayama algebras.
-The resolution quiver of $R(A)$ is the following.
-Observe that the cyclic vertex $2$ is a red vertex.
+Let $A$ be a Nakayama algebra with Kupisch series $(1,3,2,1)$. Then $A$ is the direct sum of two hereditary linear Nakayama algebras. The resolution quiver of $R(A)$ is the following.
 1234
+Observe that $4$ is a black leaf in the resolution quiver of $A$.
 For algebras of dominant dimension at least $2$, we need the following.
 Lemma 4.4.
-Let $A$ be a Nakayama algebra of dominant dimension at least $2$ and let $S$ be a simple $A$-module. If $\mathrm{pd}_{A}\,\tau^{-1}S=1$, then $\mathrm{id}_{A}\,S\neq 1$.
+Let $A$ be a Nakayama algebra. The following are equivalent.
+  1. (1)
+    $\mathrm{dom}.\dim A\geq 2$;
+  2. (2)
+    every leaf in $R(A)$ is *-black;
+  3. (3)
+    every *-red vertex in $R(A)$ is the successor of a *-black vertex.
 Proof.
-Since $\mathrm{pd}_{A}\,\tau^{-1}S=1$, then $\tau^{-1}S$ is red.
-Let $m$ be the minimal integer such that $\tau^{-m-1}S$ is black.
-By Lemma 2.1 we know that $\gamma S=\gamma\tau^{-m}S$ and $P(\tau^{-m}S)$ is injective.
-Consider the exact sequence
-$0\to P(S)\to P(\tau^{-m}S)\to I(\tau^{-1}S).$
-Since $\mathrm{dom}.\dim A\geq 2$, the module $I(\tau^{-1}S)$ is projective. Then the injective dimension of $S$ is not equal to $1$. ∎
-We have the following.
+$(1)\implies(2)$ Let $S$ be a *-red vertex in $R(A)$. Take a *-black vertex $T$ such that $\gamma(T)=\gamma(S)$. Then $P(T)$ is injective. There is an exact sequence
+$0\to P(S)\to P(T)\to I(\tau^{-1}S).$
+Since $\mathrm{dom}.\dim A\geq 2$, we have $I(\tau^{-1}S)$ is projective. Then $S$ is not a leaf.
+$(2)\implies(3)$ Let $S$ be a *-red vertex in $R(A)$. Since $S$ is not a leaf, there is a *-black vertex $T$ such that $\gamma(T)=S$. Then $S$ is the successor of a *-black vertex.
+$(3)\implies(1)$ This follows from [9, 6.2].
+Let $P(T)$ be an indecomposable non-injective projective $A$-module.
+Then $T$ is *-red and there is a *-black simple $A$-module $S$ such that $\gamma(S)=T$.
+Since $S$ is *-black, $P(S)$ and $I(\tau^{-1}T)$ are isomorphic.
+Then we have $\mathrm{dom}.\dim A\geq 2$.
+∎
 Proposition 4.5.
-Let $A$ be a Nakayama algebra and $R(A)$ be its resolution quiver. Then $A$ is minimal Auslander-Gorenstein of selfinjective dimension $2$ if and only if
+Let $A$ be a Nakayama algebra and $R(A)$ its resolution quiver. Then $A$ is minimal Auslander-Gorenstein of selfinjective dimension $2$ if and only if
   1. (1)
@@
 $\implies$” (1) If $A$ has finite global dimension, then $\operatorname{dist}A=1$ by Lemma 2.3. If $A$ has infinite global dimension, by Lemma 2.4 we have $\operatorname{dist}A=1$.
-(2) Suppose three distinct vertices $T,\tau T,\tau^{2}T$ have the same cyclic successor in $R(A)$.
+(2) Suppose three distinct vertices $T,\tau T,\tau^{2}T$ have the same successor in $R(A)$.
 Then $T$ and $\tau T$ are red.
 Since $\mathrm{dom}.\dim A\geq 2$, by Lemma 4.4 $\tau T$ and $\tau^{2}T$ are non-leaf vertices.
-By (1) we get that $\tau T,\tau^{2}T$ are cyclic.
+Since $\operatorname{dist}A=1$, both $\tau T$ and $\tau^{2}T$ are cyclic.
 This is impossible.
-(3) Let $S$ be a red vertex in $R(A)$. By (2) $\tau^{-1}S$ is black. By Lemma 2.1 $P(S)$ is injective. Since the selfinjective dimension of $A$ is $2$, there is an exact sequence
+(3) Let $S$ be a red vertex in $R(A)$. By Lemma 4.4 $\tau^{-1}S$ is black. Then $P(S)$ is injective. Since $\mathrm{id}_{A}\,A=2$, there is an exact sequence
 $0\to P(\tau S)\to P(S)\to I(S)\to I(\tau^{-1}S)\to 0.$
-Then the injective dimension of $S$ is equal to $1$. By Lemma 2.5 the vertex $S$ is a leaf in $R(A)$. In particular, $S$ is not cyclic.
+By Lemma 2.5 $S$ is a leaf in $R(A)$. In particular, $S$ is not cyclic.
 […]
-By Remark 4.2 the algebra $A$ is not hereditary.
-Then the selfinjective dimension of $A$ is $2$.
-It remains to prove $\mathrm{dom}.\dim A\geq 2$.
-By [9, 6.2], it suffices to show that each red simple $A$-module is in the image of black simple $A$-modules under $\tau^{-1}\gamma\tau$.
-Since $\operatorname{dist}A=1$ and all cyclic vertices are black.
-We infer that all black vertices are cyclic.
-Let $S$ be a red vertex in $R(A)$.
-Since $\gamma(S)$ has at most two predecessors, $\tau(S)$ is black and then cyclic.
-There is a cyclic vertex $T$ in $R(A)$ such that $\gamma(T)=\tau(S)$.
-If $\tau^{-1}T$ is black, then
-$\tau^{-1}\gamma\tau(\tau^{-1}T)=\tau^{-1}\gamma T=\tau^{-1}\tau S=S.$
-If $\tau^{-1}T$ is red, then $\tau^{-2}T$ is black. We have $\gamma\tau^{-1}T=\gamma T$ and
-$\tau^{-1}\gamma\tau(\tau^{-2}T)=\tau^{-1}\gamma\tau^{-1}T=\tau^{-1}\gamma T=\tau^{-1}\tau S=S.$
-This finishes our proof. ∎
+Let $S$ be a leaf vertex in $R(A)$.
+Since $\operatorname{dist}A=1$ and all cyclic vertices are black, then $S$ is red and $\gamma(S)=\gamma(\tau S)$.
+Since $\gamma(S)$ has at most $2$ predecessors, $\tau^{-1}S$ is black.
+By Lemma 4.4, the dominant dimension of $A$ is at least $2$.
+Then $A$ is minimal Auslander-Gorenstein of selfinjective dimension $2$.
+∎
 Example 4.6.
-Let $A$ be a Nakayama algebra with Kupisch series $(6,7,6,7)$. Then $R(A)$ is the following.
+Let $A$ be a Nakayama algebra with Kupisch series $(6,7,6,7)$. The resolution quiver $R(A)$ is the following.
 1234
-It is easy to see that $R(A)$ fulfills the condition in Proposition 4.5. Then $A$ is minimal Auslander-Gorenstein of selfinjective dimension $2$.
-For algebras of dominant dimension at least $3$, we need the following key result; see [15, 4.1 and 4.2]. We give a short proof here.
+It is easy to check that $A$ fulfills the conditions in Proposition 4.5. Then $A$ is a minimal Auslander-Gorenstein algebra of selfinjective dimension $2$.
+For algebras of dominant dimension at least $3$, recall the following key result from [16, 4]. We give a short proof here.
 Lemma 4.7.
@@
   2. (2)
-    $\tau^{-1}S$ and $S$ are black for every leaf $S$ in $R(A)$;
+    every leaf in $R(A)$ is black and *-black;
   3. (3)
@@
 Proof.
-$(1)\implies(2)$ Let $S$ be a leaf in $R(A)$.
-Since $\mathrm{dom}.\dim A\geq 2$, we know from Lemma 4.4 that $\tau^{-1}S$ is black.
-Then $P(S)$ is injective by Lemma 2.1.
-If $S$ is red, there is an exact sequence
+$(1)\implies(2)$ Let $S$ be a leaf in $R(A)$. Since $\mathrm{dom}.\dim A\geq 2$, by Lemma 4.4 $S$ is *-black and $P(S)$ is injective. Assume $S$ is red. There is an exact sequence
 $0\to P(\tau S)\to P(S)\to I(S)\to I(\tau^{-1}S)\to 0.$
-Since $\mathrm{dom}.\dim A\geq 3$, the module $I(\tau^{-1}S)$ is projective and the sequence splits. This is impossible. Then $S$ is black.
+Since $\mathrm{dom}.\dim A\geq 3$, then $I(\tau^{-1}S)$ is projective. This is impossible. We conclude that $S$ is black and *-black.
 $(2)\iff(3)$ This follows from Corollary 2.2.
-$(3)\iff(4)$ Since $R(A)^{\prime}$ and $R(\varepsilon{A})$ are isomorphic quivers, the defect of $\varepsilon(A)$ is equal to the number of leaf vertices in $R(A)^{\prime}$.
-Then $\operatorname{def}A$ is equal to $\operatorname{def}\varepsilon(A)$ if and only if $S$ is the unique predecessor of its successor for every leaf $S$ in $R(A)$.
-$(4)\implies(5)$ Let $T$ be a leaf in $R(A)$ and $S$ be the successor of $T$. Since $\tau^{-1}T$ is a black leaf, by (3.3) there is a non-split exact sequence
-$0\to J(\tau^{-1}S)\to P(T)\to P(I(\tau^{-1}T))\to I(\tau^{-1}T)\to 0.$
-Since $A$ satisfies (B), by Proposition 3.9 the projective injective objects in $\mathcal{F}$ and the projective injective $A$-modules contained in $\mathcal{F}$ coincide. Then
+$(3)\iff(4)$ Note that $\operatorname{def}\varepsilon(A)$ is equal to the number of leaf vertices in $R(A)^{\prime}$.
+Then $\operatorname{def}A$ and $\operatorname{def}\varepsilon(A)$ are equal if and only if every leaf in $R(A)$ is the unique predecessor of its successor.
+$(2)\implies(5)$ Let $T$ be a leaf in $R(A)$ and $S$ the successor of $T$. Since $\tau^{-1}T$ is black, by (3.3) there is an exact sequence
+$0\to J(\tau^{-1}S)\to P(T)\overset{w}{\to}P(I(\tau^{-1}T))\to I(\tau^{-1}T)\to 0.$
+Since $T$ is a leaf, we have $w\neq 0$. Note that $A$ satisfies (B). By Proposition 3.5 any projective injective object in $\mathcal{F}$ is a projective injective $A$-module. Then
 $\mathrm{codom}.\dim_{A}I(\tau^{-1}T)=\mathrm{codom}.\dim_{\mathcal{F}}J(\tau^{-1}S)+2.$
@@
 $(5)\implies(1)$ Since $\mathrm{dom}.\dim\varepsilon(A)\geq 1$, we have $\mathrm{dom}.\dim A\geq 3$. ∎
-Note that dominant dimension at least $3$ implies (B) and (B) for any Nakayama algebra. We have the following.
+Note that $\mathrm{dom}.\dim A\geq 3$ implies (B) and (B). We have the following.
 Corollary 4.8.
-Let $A$ be a Nakayama algebra of dominant dimension at least $3$ and let $\Delta(S)$ be a vertex in $R(\varepsilon{A})$. Then
+Let $A$ be a Nakayama algebra of dominant dimension at least $3$ and $\Delta(S)$ a vertex in $R(\varepsilon{A})$. Then
   1. (1)
@@
   2. (2)
-    $\tau^{-1}_{\varepsilon}\Delta(S)$ is black if and only if $\tau^{-1}S$ is black.
+    $\Delta(S)$ is *-black if and only if $S$ is *-black.
+Combining Lemma 4.4 and 4.7 with the previous result, we have the following.
 Corollary 4.9.
+Let $A$ be a Nakayama algebra and $m\geq 0$.
+  1. (1)
+    The following are equivalent.
+    1. (a)
+      $\mathrm{dom}.\dim A\geq 2m+1$;
+    2. (b)
+      $S,\gamma(S),\cdots,\gamma^{m-1}(S)$ are black and *-black for every leaf $S$;
+    3. (c)
+      $|\gamma^{-1}\gamma^{i}(S)|=1$ for every leaf $S$ and $i\leq m$.
+  2. (2)
+    The following are equivalent.
+    1. (a)
+      $\mathrm{dom}.\dim A\geq 2m+2$;
+    2. (b)
+      $S,\gamma(S),\cdots,\gamma^{m-1}(S)$ are black and *-black, and $\gamma^{m}(S)$ is *-black for every leaf $S$;
+    3. (c)
+      $|\gamma^{-1}\gamma^{m}(S)|=1$ and $\gamma^{m}(S)$ is *-black for every leaf $S$ and $i\leq m$.
+We have the following; compare [16, 4.4.2].
+Proposition 4.10.
 Let $A$ be a Nakayama algebra of dominant dimension $d\geq 3$.
@@
 Proof.
-This follows from Corollary 3.6 and Lemma 4.7. Compare [15, 4.4.2]. ∎
-Now we restate and prove Theorem A. It gives a criterion to decide whether a Nakayama algebra is minimal Auslander-Gorenstein of odd selfinjective dimension.
+This follows from Lemma 4.7 and Corollary 3.6. ∎
+Now we restate and prove Theorem A.
 Theorem 4.11.
-Let $A$ be a Nakayama algebra and $R(A)$ be its resolution quiver. For any $m\geq 1$, the algebra $A$ is minimal Auslander-Gorenstein of selfinjective dimension 2m12m-1 if and only if
+Let $A$ be a Nakayama algebra and $m\geq 1$. Then $A$ is minimal Auslander-Gorenstein of selfinjective dimension 2m12m-1 if and only if
   1. (1)
@@
   4. (4)
-    $\tau^{-1}$ of all cyclic vertices in $R(A)$ are black.
+    all cyclic vertices in $R(A)$ are *-black.
 Proof.
 $\implies$" We prove by induction on $m$.
-If $m=1$, this is proven in Proposition 4.1.
-If $m\geq 2$, by Corollary 4.9 the algebra $\varepsilon(A)$ is minimal Auslander-Gorenstein of selfinjective dimension 2m32m-3.
-By induction hypothesis, the following statements hold for $\varepsilon(A)$.
+If $m=1$, this is done in Proposition 4.1.
+Suppose $m\geq 2$.
+By Proposition 4.10 the syzygy filtered algebra $\varepsilon(A)$ is minimal Auslander-Gorenstein of selfinjective dimension 2m32m-3.
+By induction hypothesis, the following statements hold.
   1. (1a)
@@
   4. (4a)
-    $\tau^{-1}_{\varepsilon}$ of all cyclic vertices in $R(\varepsilon A)$ are black.
-There is an injective map of resolution quivers from $R(\varepsilon{A})$ to $R(A)$ which preserves the weight. Then (1), (2) and (3) hold. By Corollary 4.8 we get (4).
+    all cyclic vertices in $R(\varepsilon A)$ are *-black.
+Since there is an injective weight preserving map of resolution quivers from $R(\varepsilon{A})$ to $R(A)$, then (1), (2) and (3) hold. By Corollary 4.8 the statement (4) holds.
 […]
-There is an injective map of resolution quivers $\varphi$ from $R(\varepsilon{A})$ to $R(A)$ which preserves the weight. Then (1a), (2a) and (3a) hold. By Corollary 4.8 we obtain (4a).
+Since there is an injective weight preserving map of resolution quivers from $R(\varepsilon{A})$ to $R(A)$, then (1a), (2a) and (3a) hold. By Corollary 4.8 the statement (4a) holds.
 By the induction hypothesis, the algebra $\varepsilon(A)$ is minimal Auslander-Gorenstein of selfinjective dimension 2m32m-3.
-Since the dominant dimension of $A$ is at least $3$, by Corollary 4.9 the algebra $A$ is minimal Auslander-Gorenstein of selfinjective dimension 2m12m-1.
+Since the dominant dimension of $A$ is at least $3$, by Proposition 4.10 the algebra $A$ is minimal Auslander-Gorenstein of selfinjective dimension 2m12m-1.
 ∎
-We restate Theorem B as follows. The proof is similar, we omit it here.
+Let us restate Theorem B as follows. The proof is similar, we omit it here.
 Theorem 4.12.
-Let $A$ be a Nakayama algebra and $R(A)$ be its resolution quiver. For any $m\geq 1$, the algebra $A$ is minimal Auslander-Gorenstein of selfinjective dimension $2m$ if and only if
+Let $A$ be a Nakayama algebra and $m\geq 1$. Then $A$ is minimal Auslander-Gorenstein of selfinjective dimension $2m$ if and only if
   1. (1)
@@
     all cyclic vertices in $R(A)$ are black.
-For a Nakayama algebra $A$, denote by $\operatorname{cyc}A$ be the number of cyclic vertices in its resolution quiver. We have the following.
+For a Nakayama algebra $A$, denote by $\operatorname{cyc}A$ the number of cyclic vertices in its resolution quiver.
 Corollary 4.13.
-Let $A$ be a Nakayama algebra. Then $\operatorname{def}A\leq\operatorname{cyc}A$ provided that $A$ is minimal Auslander-Gorenstein of even selfinjective dimension.
+Let $A$ be a Nakayama algebra. If $A$ is a minimal Auslander-Gorenstein algebra of even selfinjective dimension, then $\operatorname{cyc}A\leq\operatorname{def}A$.
 5. Applications and examples
-Let $A$ be a Nakayama algebra of rank $n$ with Kupisch series
+In this section, we present some applications and examples of our results.
+Let $A$ be a Nakayama algebra with Kupisch series
 $(c_{1},c_{2},\cdots,c_{n}).$
@@
 Proposition 5.1.
-Let $A$ be a Nakayama algebra and let $d$ be an even number. Then $A$ is minimal Auslander-Gorenstein of selfinjective dimension $d$ if and only if so is $\bar{A}$.
-For any integer $n$, denote by $\mathbb{A}_{n}$ be the hereditary linear Nakayama algebra of rank $n$.
-For any $m\leq n$, denote by $N_{n}(m)$ be the quotient algebra of $\mathbb{A}_{n}$ via the ideal generated by all paths of length $m$.
+Let $A$ be a Nakayama algebra and $d$ an even number. Then $A$ is minimal Auslander-Gorenstein of selfinjective dimension $d$ if and only if so is $\bar{A}$.
+Let $2\leq m\leq n$.
+Denote by $\mathbb{A}_{n}$ the hereditary linear Nakayama algebra of rank $n$.
+Let $N_{n}(m)$ be the quotient algebra of $\mathbb{A}_{n}$ via the ideal generated by all paths of length $m$.
+The algebra $N_{n}(m)$ is a linear Nakayama algebra and has finite global dimension.
 The Kupisch series of $N_{n}(m)$ is
 $((m)^{n-m+1},m-1,\cdots,2,1).$
-The algebra $N_{n}(m)$ has finite global dimension. If $m=2$, then $N_{n}(m)$ is higher Auslander of global dimension n1n-1.
-Suppose $m>2$ and $N_{n}(m)$ is a higher Auslander algebra. Write
+If $m=2$, then $N_{n}(m)$ is higher Auslander of global dimension n1n-1. Suppose $m>2$ and $N_{n}(m)$ is a higher Auslander algebra. We write
 $n=qm+r$
 such that $q,r$ are integers and $0\leq r\leq m-1$.
-Since $S(1)$ is cyclic and $\gamma^{-1}S(1)$ consists of $m$ simple $A$-modules, the global dimension of $N(n,r)$ is odd by Theorem 4.11.
-Since S(nr+1)S(n-r+1) is cyclic, then S(nr)S(n-r) is black by Theorem 4.10.
-We infer that $r=0$ and $m$ divides $n$.
+Since $S(1)$ is cyclic and $S(1)$ has $m$ predecessors in the resolution quiver, the global dimension of $N_{n}(m)$ is odd by Theorem 4.12.
+Since S(nr+1)S(n-r+1) is cyclic, S(nr)S(n-r) is black by Theorem 4.11.
+It follows that $r=0$ and $m$ divides $n$.
 If $m$ divides $n$, it is routine to check that $N_{n}(m)$ is higher Auslander of global dimension 2n/m12n/m-1.
 Proposition 5.2.
-The algebra $N_{n}(m)$ is higher Auslander if and only if $m=2$ or $m$ divides $n$. In both cases, the global dimension of $N_{n}(m)$ is 2n/m12n/m-1.
-Following [15], for any Nakayama algebra $A$, there is a unique Nakayama algebra $\varepsilon^{-1}{A}$ of dominant dimension at least $3$ such that $\varepsilon(\varepsilon^{-1}{A})$ is isomorphic to $A$.
-Let $A$ be a Nakayama algebra of rank $n$ with Kupisch series
+The algebra $N_{n}(m)$ is a higher Auslander algebra if and only if $m=2$ or $m$ divides $n$. In both cases, the global dimension of $N_{n}(m)$ is 2n/m12n/m-1.
+For any Nakayama algebra $A$, by [16] there is a unique Nakayama algebra $\varepsilon^{-1}{A}$ of dominant dimension $\geq 3$ such that $\varepsilon(\varepsilon^{-1}{A})$ is naturally isomorphic to $A$.
+The algebra $\varepsilon^{-1}A$ is given as follows.
+Suppose the Kupisch series of $A$ is
 $c(A)=(c_{1},c_{2},\cdots,c_{n}).$
-Denote by $d_{a}$ the maximum of $c_{a}-c_{a-1}$ and 0 for each $a\leq n$, and denote by $d$ the sum of all $d_{a}$. Let $\tilde{c}_{a}$ be a series of length $1+d_{a+1}$ such that
-$(\tilde{c}_{a})_{\ell}=c_{a}+\sum_{i=1}^{c_{a}+\ell-1}d_{a+i}$
-for each $\ell\leq 1+d_{a+1}$. Then $\varepsilon^{-1}{A}$ is a Nakayama algebra of rank $n+d$ with Kupisch series
+Denote by $d_{a}$ the maximum of $c_{a}-c_{a-1}$ and 0 for each $a\leq n$, and by $d$ the sum of all $d_{a}$.
+The subscript is taken modulo $n$.
+Let $\tilde{c}_{a}$ be a series of length $1+d_{a+1}$ with
+[…]
+Then $\varepsilon^{-1}{A}$ is a Nakayama algebra of rank $n+d$ with Kupisch series
 $c(\varepsilon^{-1}A)=(\tilde{c}_{1},\tilde{c}_{2},\cdots,\tilde{c}_{n}).$
-One sees that $\chi(\varepsilon^{-1}{A})$ is equal to the number of simple components of $A$.
-Recall the following result in [15].
+Note that $\chi(\varepsilon^{-1}A)$ equals the number of simple components of $A$.
+Recall the following result; see [14, 16].
 Proposition 5.3.
-Let $d\geq 3$ and $\Lambda$ be a cyclic Nakayama algebra. Then $\Lambda$ is higher Auslander of global dimension $d$ if and only if $\Lambda$ is isomorphic to
+Let $\Lambda$ be a cyclic Nakayama algebra and $d\geq 3$. Then $\Lambda$ is higher Auslander of global dimension $d$ if and only if $\Lambda$ is isomorphic to
 $\varepsilon^{-1}(A)$
-where $A$ is a higher Auslander cyclic Nakayama algebra of global dimension d2d-2 or
+where $A$ is a cyclic Nakayama algebra which is higher Auslander of global dimension d2d-2 or isomorphic to
 $\varepsilon^{-1}(A_{1}\oplus A_{2}\oplus\cdots\oplus A_{r})$
-where $A_{i}$ is a higher Auslander linear Nakayama algebra of global dimension d2d-2 for $i\leq r$.
-Similarly, we have the following.
+where $A_{i}$ is a linear Nakayama algebra which is higher Auslander of global dimension d2d-2 for each $i\leq r$.
+By Proposition 4.10, we have the following.
 Proposition 5.4.
-Let $d\geq 3$ and $\Lambda$ be a linear Nakayama algebra. Then $\Lambda$ is higher Auslander of global dimension $d$ if and only if $\Lambda$ is isomorphic to
+Let $\Lambda$ be a linear Nakayama algebra and $d\geq 3$. Then $\Lambda$ is higher Auslander of global dimension $d$ if and only if $\Lambda$ is isomorphic to
 $\varepsilon^{-1}(\mathbb{A}_{1}\oplus A_{1}\oplus\cdots\oplus A_{r})$
-where $A_{i}$ is a higher Auslander linear Nakayama algebra of global dimension d2d-2 for $i\leq r$.
-Connected Nakayama algebras of which are higher Auslander of global dimension $1$ are hereditary.
-Connected Nakayama algebras of which are higher Auslander of global dimension $2$ are just the Auslander algebras of radical square zero connected Nakayama algebras.
-We can repeatedly apply Proposition 5.4 and 5.3 to give all connected Nakayama algebras of which are higher Auslander.
+where $A_{i}$ is a linear Nakayama algebra which is higher Auslander of global dimension d2d-2 for each $i\leq r$.
 Example 5.5.
@@
 $(1,3,2,1,2,1).$
-The Kupisch series of $\varepsilon^{-1}(A)$ is
+Then $\varepsilon^{-1}(A)$ is higher Auslander of global dimension $3$ with Kupisch series
 $(3,3,3,4,3,2,2,2,1).$
-Then $\varepsilon^{-1}(A)$ is higher Auslander of global dimension $3$.
+Nakayama algebras of global dimension $1$ are hereditary, and Nakayama algebras which are higher Auslander of global dimension $2$ are the Auslander algebras of Nakayama algebras with radical square zero.
+For any $d\geq 3$, we can repeatedly apply Proposition 5.4 and 5.3 to get any connected Nakayama algebra which is higher Auslander of global dimension $d$.
+Example 5.6.
+Let $n\geq m\geq 0$ and $A$ be a non-simple linear Nakayama algebra. The algebra
+$\Lambda=\varepsilon^{-n}(\mathbb{A}_{1}^{m}\oplus A)$
+is a connected Nakayama algebra; it is linear if and only if $n=m$. We have
+  1. (1)
+    $\varepsilon^{-m}(\mathbb{A}_{1}^{m}\oplus\mathbb{A}_{n})\cong N_{mn+n}(n)$ for any $m\geq 1$ and $n\geq 2$;
+  2. (2)
+    $\varepsilon^{-m}(\mathbb{A}_{1}^{m}\oplus N_{3}(2))\cong N_{2m+3}(2)$ for any $m\geq 1$.